### v1

# TV homogenization inequalities

Abstract
We study the total variation distance under two information-erasing maps on inhomogeneous Bernoulli product measures: summation and homogenization. While summation is a Markov kernel and hence satisfies the usual data processing inequality, homogenization — which maps each Bernoulli parameter to the cumulative mean — is not. Nevertheless, we prove that the homogenization map also reduces the TV distance, up to a universal constant. The argument is based on an explicit two-sided control of the TV distance between Poisson binomials, obtained via a parameter interpolation and a second-moment extraction.
1 Introduction
For $\textup{{\bf p}}=(p_{1},\dots,p_{n})\in[0,1]^{n}$, the Poisson–binomial law, introduced by [10] [^1] See [11] for a modern English translation. , is the pushforward of the product measure $\operatorname{Ber}(\textup{{\bf p}}):=\operatorname{Ber}(p_{1})\otimes\cdots\otimes\operatorname{Ber}(p_{n})$ under the sum map $x\mapsto\sum_{i}x_{i}$. Equivalently, if $X_{i}\sim\operatorname{Ber}(p_{i})$ are independent, then $S_{\textup{{\bf p}}}:=\sum_{i=1}^{n}X_{i}$ is distributed according to this law. The sum map is easily seen to be a Markov kernel, and hence the Data Processing Inequality [12, Theorem 7.4] applies: $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$, where $\operatorname{TV}$ is the total variation distance. Intuitively, since information is lost under the sum map, the distributions can only become closer to each other.
In this paper, we investigate the behavior of TV under the sum map at the parameter rather than the observation level. The homogenization map replaces each $p_{i}$ with $\bar{p}=\frac{1}{n}\sum_{i}p_{i}$, transforming $\operatorname{Ber}(\textup{{\bf p}})$ into $\operatorname{Ber}(\bar{p})^{\otimes n}$. Since this operation also erases information, we might likewise expect it to decrease the TV distance. One immediate glitch is that the homogenization map cannot be realized by any Markov kernel. Indeed, suppose to the contrary that for $n\geq 2$, some Markov kernel $K$ maps $X\sim\operatorname{Ber}(\textup{{\bf p}})$ to $Y\sim\operatorname{Ber}(\bar{p})^{\otimes n}$. Applying the sum map to $Y$ yields $Z\sim\operatorname{Bin}(n,\bar{p})$; thus we have some Markov kernel $\tilde{K}:\operatorname{Ber}(\textup{{\bf p}})\mapsto\operatorname{Bin}(n,\bar{p})$ for all $\textup{{\bf p}}\in[0,1]^{n}$. Consider a p of the form $\textup{{\bf p}}=(t,0,0,\ldots,0)\in[0,1]^{n}$. Then the law of $X$ is affine in $t$ (it is a $t$-convex combination of two atoms), as is the pushforward measure under $\tilde{K}$. However, $\mathbb{P}(Z=2)=\binom{n}{2}\left(\frac{t}{n}\right)^{2}\left(1-\frac{t}{n}\right)^{n-2}$, which is not affine in $t$.
This indicates that the standard information-theoretic argument for proving data processing inequalities does not apply, necessitating novel tools. One such tool we propose analyzes the behavior of the Poisson binomial $S_{\textup{{\bf p}}}$ under the total variation metric $\operatorname{TV}$. For $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, define the quantities
$\Delta:=\sum_{i=1}^{n}|p_{i}-q_{i}|,\quad\sigma_{\textup{{\bf p}}}^{2}:=\operatorname{Var}(S_{\textup{{\bf p}}})=\sum_{i=1}^{n}p_{i}(1-p_{i}),\quad\Phi(\textup{{\bf p}},\textup{{\bf q}}):=\min\!\left(1,\frac{\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}}\right).$ (1)
We discover a simple analytic approximation of $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ for dominating pairs $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, $p_{i}\geq q_{i}$, up to universal multiplicative constants:
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\asymp\Phi(\textup{{\bf p}},\textup{{\bf q}}),\qquad\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n},\textup{{\bf p}}\geq\textup{{\bf q}}.$ (2)
For general $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, only the upper bound $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\lesssim\Phi(\textup{{\bf p}},\textup{{\bf q}})$ holds, which can be sharpened to $\min(\Phi(\textup{{\bf p}},\textup{{\bf q}}),\Phi(\textup{{\bf q}},\textup{{\bf p}}))$ by symmetry. Unlike $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$, whose exact computation is infeasible (see below), $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ is computable exactly in time $O(n^{2})$ via a simple recurrence [15, Section 5]. Still, one should not expect any analytically tractable analogue of (2) for general $\textup{{\bf p}},\textup{{\bf q}}$ due to cancellations and convolutional smoothing.
As an application of these results, we prove the TV homogenization inequality for all $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$:
$\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\gtrsim$ $\operatorname{TV}(\operatorname{Ber}(\bar{p})^{\otimes n},\operatorname{Ber}(\bar{q})^{\otimes n})=\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q})).$ (3)
Notation.
We write $n]:=\left\{1,\ldots,n\right\}$. For two distributions $P,Q$ on a finite set $\Omega$, their total variation distance is defined by
$\operatorname{TV}(P,Q)$ $=$ $\frac{1}{2}\sum_{\omega\in\Omega}|P(\omega)-Q(\omega)|.$ (4)
If $X,Y$ are random variables with laws $P,Q$, respectively, then we write $\operatorname{TV}(X,Y):=\operatorname{TV}(P,Q)$. For $p\in[0,1]$, $\operatorname{Ber}(p)$ denotes the Bernoulli measure on $\left\{0,1\right\}$: $\operatorname{Ber}(p)(0)=1-p$ and $\operatorname{Ber}(p)(1)=p$. For $n\in\mathbb{N}$ and $\textup{{\bf p}}\in[0,1]^{n}$, $\operatorname{Ber}(\textup{{\bf p}})$ denotes the product of $n$ Bernoulli distributions with parameters $p_{i}$: $\operatorname{Ber}(\textup{{\bf p}})=\operatorname{Ber}(p_{1})\otimes\operatorname{Ber}(p_{2})\otimes\ldots\otimes\operatorname{Ber}(p_{n}).$ The Poisson binomial $S_{\textup{{\bf p}}}$ is a sum of $n$ independent $\operatorname{Ber}(p_{i})$ variables. When all of the $p_{i}$ are identical (say, to $\bar{p}$), this is the binomial distribution $\operatorname{Bin}(n,\bar{p})$. The notation in (1) will persist throughout the paper.
Main results.
Our first result implies the upper bound in (2). Let $\eta_{\mathrm{BCV}}$ denote the sharp universal constant from [1, Theorem 1], and set
$C_{\mathrm{BCV}}:=\sqrt{\frac{5}{4}+\eta_{\mathrm{BCV}}^{2}}.$
Numerically (see [1, Remark 1]), $\eta_{\mathrm{BCV}}\approx 0.4688223555$ and hence $C_{\mathrm{BCV}}\approx 1.2123507747$.
Theorem 1.
For $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ and $\Phi$ as in (1),
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\leq$ $C_{\mathrm{BCV}}\,\min(\Phi(\textup{{\bf p}},\textup{{\bf q}}),\Phi(\textup{{\bf q}},\textup{{\bf p}})).$
The technical core of the paper is a matching lower bound for dominating pairs:
Theorem 2.
For $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ satisfy $p_{i}\geq q_{i}$, $i\in[n]$, and $\Phi$ as in (1),
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\geq$ $\frac{1}{12}\Phi(\textup{{\bf p}},\textup{{\bf q}}).$
The assumption $\textup{{\bf p}}\geq\textup{{\bf q}}$ is essential for the lower bound, as illustrated by the example $\textup{{\bf p}}=(1,\,0,\,\tfrac{1}{2})$ and $\textup{{\bf q}}=(0,\,1,\,\tfrac{1}{2}+\varepsilon)$; here, $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})=\varepsilon$ while $\Phi(\textup{{\bf p}},\textup{{\bf q}})=1$. In general, for any $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ and any partition $I,J$ of $[n]$, the trivial estimate via the partitioned pairs $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\leq\operatorname{TV}(S_{\textup{{\bf p}}_{I}},S_{\textup{{\bf q}}_{I}})+\operatorname{TV}(S_{\textup{{\bf p}}_{J}},S_{\textup{{\bf q}}_{J}})$ holds. However, this method cannot yield a matching lower bound, as illustrated by the example $a,b=\tfrac{1}{2}\pm\varepsilon$, $\textup{{\bf p}}=(a,a,b,b)$, $\textup{{\bf q}}=(a,b,a,b)$ with the partition $I=\left\{1,2\right\}$, $J=\left\{3,4\right\}$.
Although for general $\textup{{\bf p}},\textup{{\bf q}}$, the approach of Theorem 2 does not lower-bound $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$, it does yield a lower bound for the larger quantity $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$:
Theorem 3.
For $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, let $I=\left\{i\in[n]:p_{i}\geq q_{i}\right\}$ and $J=[n]\setminus I$, with $\textup{{\bf p}}_{I},\textup{{\bf p}}_{J},\textup{{\bf q}}_{I},\textup{{\bf q}}_{J}$ the corresponding subsequences of $\textup{{\bf p}},\textup{{\bf q}}$. Then
$\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\geq$ $\max\left(\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}_{I}),\operatorname{Ber}(\textup{{\bf q}}_{I})),\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}_{J}),\operatorname{Ber}(\textup{{\bf q}}_{J}))\right)$
$\geq$ $\frac{1}{12}\max(\Phi(\textup{{\bf p}}_{I},\textup{{\bf q}}_{I}),\Phi(\textup{{\bf q}}_{J},\textup{{\bf p}}_{J})).$
The last piece to complete our program is a homogenization result for binomials:
Lemma 4.
For $n\geq 1$, $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, and $\emptyset\neq A\subset N:=[n]$, define $\bar{p}_{A}=|A|^{-1}\sum_{i\in A}p_{i}$ and $\bar{q}_{A}$ analogously; put
$\delta_{A}$ $:=$ $\operatorname{TV}(\operatorname{Bin}(|A|,\bar{p}_{A}),\operatorname{Bin}(|A|,\bar{q}_{A})).$
If $I,J$ form a partition of $N$ with $I,J\neq\emptyset$, then
$\delta_{N}\;\leq\;2\bigl(\delta_{I}+\delta_{J}\bigr).$
The homogenization inequality claimed in (3) then follows:
Theorem 5.
For all $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, we have
$\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\geq$ $c\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q})),$
where $c\geq\frac{1}{48C_{\mathrm{BCV}}}\approx 0.017$ is a universal constant.
Remark.
The choice $\textup{{\bf p}}=(1-2\varepsilon,\tfrac{1}{2})$, $\textup{{\bf q}}=(1,\tfrac{1}{2}+\varepsilon)$, $\varepsilon\to 0$ demonstrates that it is possible for homogenization to (slightly) increase the TV distance, and also that $8/9$ is an upper bound on $c$. We conjecture that this value is in fact optimal, see below.
Proof.
Partition $[n]$ into $I,J$ as in Theorem 3 and ignore the trivial case where one of $I,J$ is empty. Then Theorem 3 implies
$\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\geq$ $\frac{1}{12}\max(\Phi(\textup{{\bf p}}_{I},\textup{{\bf q}}_{I}),\Phi(\textup{{\bf q}}_{J},\textup{{\bf p}}_{J}))$
$\geq$ $\frac{1}{24}\left(\Phi(\textup{{\bf p}}_{I},\textup{{\bf q}}_{I})+\Phi(\textup{{\bf q}}_{J},\textup{{\bf p}}_{J})\right).$
A simple convexity argument shows that homogenization cannot increase $\Phi$:
$\Delta(\textup{{\bf p}},\textup{{\bf q}})$ $\geq$ $\Delta(\bar{p}\boldsymbol{1},\bar{q}\boldsymbol{1}),$
$\sigma_{\textup{{\bf p}}}^{2}$ $\leq$ $\sigma_{\bar{p}\boldsymbol{1}}^{2}=n\bar{p}(1-\bar{p}),$
where $\bar{p}\boldsymbol{1}=(\bar{p},\bar{p},\ldots,\bar{p})$ is of the same dimension as p.
Now homogenize $\bar{p}_{I}$, $\bar{p}_{J}$, $\bar{q}_{I}$, $\bar{q}_{J}$ as in Lemma 4. Since $\operatorname{Bin}(|I|,\bar{p}_{I})$ is a special case of Poisson binomial, Theorem 1 applies:
$\Phi(\textup{{\bf p}}_{I},\textup{{\bf q}}_{I})+\Phi(\textup{{\bf q}}_{J},\textup{{\bf p}}_{J})$ $\geq$ $\Phi(\bar{p}_{I}\boldsymbol{1},\bar{q}_{I}\boldsymbol{1})+\Phi(\bar{q}_{J}\boldsymbol{1},\bar{p}_{J}\boldsymbol{1})$
$\geq$ $\frac{1}{C_{\mathrm{BCV}}}\left(\operatorname{TV}(\operatorname{Bin}(|I|,\bar{p}_{I}),\operatorname{Bin}(|I|,\bar{q}_{I}))+\operatorname{TV}(\operatorname{Bin}(|J|,\bar{p}_{J}),\operatorname{Bin}(|J|,\bar{q}_{J}))\right)$
$\geq$ $\frac{1}{2C_{\mathrm{BCV}}}\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q})),$
where the last inequality is by Lemma 4. Finally, the identity $\operatorname{TV}(\operatorname{Ber}(\bar{p})^{\otimes n},\operatorname{Ber}(\bar{q})^{\otimes n})$ == $\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q}))$ is a standard consequence of the Neyman–Pearson lemma (the likelihood ratio is determined by and monotone in the sum). ∎
Remark and open problems.
Although the homogenization inequality relies on novel structural insights into the Poisson binomial, there are compelling reasons to believe that a great deal more structure remains to be uncovered, currently out of reach. Indeed, copious numerical trials unambiguously indicate that the correct constant in Theorem 5 should be $c=8/9$ and also that the bound in Lemma 4 can be sharpened to $\delta_{N}\leq\delta_{I}+\delta_{J}-\delta_{I}\delta_{J}$ (the latter, in particular, appears deceptively simple, since only binomials are involved). Our present methods do not seem to provide any pathway to these conjecturally optimal bounds, which will have to await further structural advances.
Related work.
This paper follows a program initiated by [8] to provide simple, analytically tractable upper and lower estimates on $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ in terms of the $\textup{{\bf p}},\textup{{\bf q}}$; the latter’s contribution was a sharpening of the trivial lower bound $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{\infty}$ to $\gtrsim\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{2}$. We note that generally, the latter is incomparable with the lower bound $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\gtrsim\Phi(\textup{{\bf p}},\textup{{\bf q}})$ implied by Theorem 2. Indeed, the choice $q_{i}\equiv\frac{1}{2}$ for $i\in[n]$ and $p_{i}=1/2$, $i>1$ and $p_{1}=\frac{1}{2}+\varepsilon$ yields $\Phi(\textup{{\bf p}},\textup{{\bf q}})\asymp\varepsilon/\sqrt{n}$ while $\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{2}=\varepsilon$. For the other direction, the choice $\textup{{\bf q}}\equiv 0$ and $\textup{{\bf p}}\equiv 1/n$ yields $\Phi(\textup{{\bf p}},\textup{{\bf q}})\asymp 1$ and $\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|=n^{-1/2}$. Homogenization inequalities appear not to have been widely studied; one classic result due to Hoeffding [7] is that homogenization under a fixed-mean constraint maximizes $\mathbb{E}g(S_{\textup{{\bf p}}})$ for any convex $g$.
The study of $S_{\textup{{\bf p}}}$ is amply motivated by numerous applications and a venerable line of work has considered approximating this distribution by simpler simpler ones [15]; perhaps most famous is Le Cam’s inequality $\operatorname{TV}(S_{\textup{{\bf p}}},\mathrm{Poi}(\sum p_{i}))\leq\sum{p_{i}^{2}}$ [9]. Note, however, that the latter yields an additive, rather than multiplicative approximation to $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$. Binomial approximation bounds in strong metrics (including $\operatorname{TV}$) appear in [4] and refinements based on orthogonal-polynomial expansions and asymptotics in [13]. Later, [14] obtained TV bounds for approximating general convolutions by compound Poisson laws using Kerstan’s method together with novel “smoothness inequalities”. A delicate structural result due to [3] (also ultimately yielding an additive estimate) shows that every Poisson binomial admits an $\varepsilon$-approximation in TV by a distribution of one of two canonical “compressed” types: either a sparse $S_{\textup{{\bf p}}}$ with only $O(1/\varepsilon^{3})$ nontrivial summands, or a near-binomial form. Of particular relevance to this work is [1], whose structural results we build upon in Theorem 1.
More generally, $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ admits classic approximations in terms of more analytically tractable proxies such as KL-divergence and the Hellinger distance; their properties and limitations are discussed in [8]. Regarding the algorithmic aspect, [2] showed that computing $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ exactly for general $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ is hard in the #\#P sense. An efficient randomized algorithm guaranteeing a $1\pm\varepsilon$ multiplicative approximation with confidence $1-\delta$, in time $O(\frac{n^{2}}{\varepsilon^{2}}\log\frac{1}{\delta})$ was discovered by [6], and later derandomized by [5].
2 Proofs
2.1 Proof of Theorem 1
We will prove a sharper, symmetric bound, which immediately implies the one in Theorem 1:
Theorem 6.
For $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ and $S_{\textup{{\bf p}}},S_{\textup{{\bf q}}}$ the corresponding Poisson binomials with variances $\sigma_{\textup{{\bf p}}}^{2}$, $\sigma_{\textup{{\bf q}}}^{2}$, we have
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\leq$ $\frac{2C_{\mathrm{BCV}}\,\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+\sqrt{\sigma_{\textup{{\bf q}}}^{2}+1}}.$
Proof.
The proof uses three standard ingredients.
Shift-TV for unimodal pmfs. If $Z$ is integer-valued with unimodal pmf $h(k)=\mathbb{P}(Z=k)$, then
$\operatorname{TV}(Z,Z+1)=\max_{k}h(k),$ (5)
by a simple telescoping argument.
Anti-concentration for Poisson binomials (Baillon–Cominetti–Vaisman). If $Z=\sum_{j=1}^{m}\operatorname{Ber}(r_{j})$ has variance $v=\operatorname{Var}(Z)$, then [1, Theorem 1] proves the sharp bound
$\max_{k}\mathbb{P}(Z=k)\leq\frac{\eta_{\mathrm{BCV}}}{\sqrt{v}}\qquad(v>0),$ (6)
where $\eta_{\mathrm{BCV}}$ is a universal constant ($\eta_{\mathrm{BCV}}\approx 0.4688223555$, [1, Remark 1]). In particular, since $\max_{k}\mathbb{P}(Z=k)\leq 1$ always, we have for all $v\geq 0$
$\max_{k}\mathbb{P}(Z=k)\leq\min\!\left\{1,\ \frac{\eta_{\mathrm{BCV}}}{\sqrt{v}}\right\}.$ (7)
Unimodality of Poisson binomial pmfs. It is a classic consequence of the Aissen-Endrei-Schoenberg-Whitney theorem [15, Section 4] that the Poisson binomial has a log-concave, and hence unimodal, law.
To combine these three ingredients, we begin with the interpolation. For $t\in[0,1]$ define
$r_{i}(t):=(1-t)q_{i}+tp_{i},\qquad S(t):=\sum_{i=1}^{n}\operatorname{Ber}(r_{i}(t)),\qquad t\in[0,1],$
where the Bernoullis are independent; thus, $S(0)=S_{\textup{{\bf q}}}$ and $S(1)=S_{\textup{{\bf p}}}$ in distribution. For $A\subseteq\mathbb{Z}$, we write $f_{A}(t):=\mathbb{P}(S(t)\in A)$. For each $i\in[n]$ let $X_{i}(t)\sim\operatorname{Ber}(r_{i}(t))$ denote the $i$th summand and set
$T_{i}(t):=S(t)-X_{i}(t)=\sum_{j\neq i}\operatorname{Ber}(r_{j}(t)).$
Then $T_{i}(t)$ is independent of $X_{i}(t)$ and $S(t)=T_{i}(t)+X_{i}(t)$.
Conditioning on $X_{i}(t)$ gives
$f_{A}(t)=r_{i}(t)\mathbb{P}(T_{i}(t)+1\in A)+(1-r_{i}(t))\mathbb{P}(T_{i}(t)\in A).$
View $f_{A}$ as a multilinear polynomial in the coordinates $(r_{1},\dots,r_{n})$. By the chain rule, $\frac{\mathrm{d}}{\mathrm{d}t}f_{A}(r(t))=\sum_{i=1}^{n}r_{i}^{\prime}(t)\,\partial f_{A}/\partial r_{i}$, and conditioning on $X_{i}(t)$ yields $\partial f_{A}/\partial r_{i}=\mathbb{P}(T_{i}(t)+1\in A)-\mathbb{P}(T_{i}(t)\in A)$. Differentiating and using $r_{i}^{\prime}(t)=p_{i}-q_{i}$ yields
$f_{A}^{\prime}(t)=\sum_{i=1}^{n}(p_{i}-q_{i})\Bigl(\mathbb{P}(T_{i}(t)+1\in A)-\mathbb{P}(T_{i}(t)\in A)\Bigr).$
Hence, by the definition of total variation,
$|f_{A}^{\prime}(t)|\leq\sum_{i=1}^{n}|p_{i}-q_{i}|\,\operatorname{TV}\!\bigl(T_{i}(t)+1,T_{i}(t)\bigr).$
Each $T_{i}(t)$ is Poisson binomial, hence unimodal, so by (5),
$\operatorname{TV}\!\bigl(T_{i}(t)+1,T_{i}(t)\bigr)=\max_{k}\mathbb{P}(T_{i}(t)=k).$
By (7),
$\max_{k}\mathbb{P}(T_{i}(t)=k)\leq\min\!\left\{1,\ \frac{\eta_{\mathrm{BCV}}}{\sqrt{\operatorname{Var}(T_{i}(t))}}\right\}.$
Since $S(t)=T_{i}(t)+X_{i}(t)$ with independence and $\operatorname{Var}(X_{i}(t))\leq\tfrac{1}{4}$, we have
$\operatorname{Var}(T_{i}(t))=\operatorname{Var}(S(t))-\operatorname{Var}(X_{i}(t))\geq\operatorname{Var}(S(t))-\frac{1}{4}.$
Let $x_{+}:=\max\{x,0\}$. Then
$\max_{k}\mathbb{P}(T_{i}(t)=k)\leq\min\!\left\{1,\ \frac{\eta_{\mathrm{BCV}}}{\sqrt{\bigl(\operatorname{Var}(S(t))-\frac{1}{4}\bigr)_{+}}}\right\}.$
We now upper bound this “singular” expression by a smooth envelope. Set $C_{\mathrm{BCV}}:=\sqrt{\tfrac{5}{4}+\eta_{\mathrm{BCV}}^{2}}$. Then for every $x\geq 0$ we have
$\min\!\left\{1,\ \frac{\eta_{\mathrm{BCV}}}{\sqrt{(x-\frac{1}{4})_{+}}}\right\}\leq\frac{C_{\mathrm{BCV}}}{\sqrt{x+1}}.$ (8)
Indeed, if $x\leq\tfrac{1}{4}+\eta_{\mathrm{BCV}}^{2}$, the left-hand side equals $1$ and the right-hand side is $C_{\mathrm{BCV}}/\sqrt{x+1}\geq C_{\mathrm{BCV}}/\sqrt{\tfrac{5}{4}+\eta_{\mathrm{BCV}}^{2}}=1$. If $x\geq\tfrac{1}{4}+\eta_{\mathrm{BCV}}^{2}$, then the left-hand side equals $\eta_{\mathrm{BCV}}/\sqrt{x-\tfrac{1}{4}}$ and squaring shows that (8) is equivalent to $x\geq\tfrac{1}{4}+\eta_{\mathrm{BCV}}^{2}$.
Applying (8) with $x=\operatorname{Var}(S(t))$ gives
$\operatorname{TV}\!\bigl(T_{i}(t)+1,T_{i}(t)\bigr)\leq\frac{C_{\mathrm{BCV}}}{\sqrt{\operatorname{Var}(S(t))+1}},$
and substituting back,
$|f_{A}^{\prime}(t)|\leq\frac{C_{\mathrm{BCV}}\,\Delta}{\sqrt{\operatorname{Var}(S(t))+1}}.$ (9)
Since $u\mapsto u(1-u)$ is concave,
$r_{i}(t)(1-r_{i}(t))\geq(1-t)\,q_{i}(1-q_{i})+t\,p_{i}(1-p_{i}),\qquad i\in[n],$
and summing over $i$ yields
$\operatorname{Var}(S(t))\geq(1-t)\sigma_{\textup{{\bf q}}}^{2}+t\sigma_{\textup{{\bf p}}}^{2}.$
Combining with (9),
$|f_{A}^{\prime}(t)|\leq\frac{C_{\mathrm{BCV}}\,\Delta}{\sqrt{(1-t)\sigma_{\textup{{\bf q}}}^{2}+t\sigma_{\textup{{\bf p}}}^{2}+1}}.$
Integrating over $t$ gives
$\bigl|f_{A}(1)-f_{A}(0)\bigr|\leq C_{\mathrm{BCV}}\,\Delta\int_{0}^{1}\frac{\mathrm{d}t}{\sqrt{(1-t)(\sigma_{\textup{{\bf q}}}^{2}+1)+t(\sigma_{\textup{{\bf p}}}^{2}+1)}}.$
Write $(1-t)(\sigma_{\textup{{\bf q}}}^{2}+1)+t(\sigma_{\textup{{\bf p}}}^{2}+1)=b+at$ with $b=\sigma_{\textup{{\bf q}}}^{2}+1$ and $a=\sigma_{\textup{{\bf p}}}^{2}-\sigma_{\textup{{\bf q}}}^{2}$. If $a=0$, the integral is $1/\sqrt{b}$. If $a\neq 0$, then
$\int_{0}^{1}\frac{\mathrm{d}t}{\sqrt{b+at}}=\left[\frac{2}{a}\sqrt{b+at}\right]_{0}^{1}=\frac{2}{\sqrt{b+a}+\sqrt{b}}=\frac{2}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+\sqrt{\sigma_{\textup{{\bf q}}}^{2}+1}}.$
Thus, for every event $A$,
$\bigl|\mathbb{P}(S_{\textup{{\bf p}}}\in A)-\mathbb{P}(S_{\textup{{\bf q}}}\in A)\bigr|\leq\frac{2C_{\mathrm{BCV}}\,\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+\sqrt{\sigma_{\textup{{\bf q}}}^{2}+1}}.$
Taking $\sup_{A}$ proves (6). ∎
2.2 Proof of Theorem 2
We begin with an abstract result showing that if a positive measure $g$ on $\mathbb{Z}$ has total mass $G$ and its variance-like spread about $\mu$ is controlled by $J<\infty$, this implies a lower bound on the mass of the heaviest atom.
Lemma 7 (Pigeonhole).
For $g:\mathbb{Z}\to[0,\infty)$ and $\mu\in\mathbb{R}$, define
$G:=\sum_{k\in\mathbb{Z}}g(k)\qquad J:=\sum_{k\in\mathbb{Z}}(k-\mu)^{2}\,g(k).$
If $0<G,J<\infty$, then
$\sup_{k\in\mathbb{Z}}g(k)\;\geq\;\frac{3\,G^{3/2}}{16\sqrt{J}+4\sqrt{G}}.$
Proof.
Put $R:=2\sqrt{{J}/{G}}$ and define
$A:=\{k\in\mathbb{Z}:\ |k-\mu|\leq R\},\qquad A^{\mathrm{c}}:=\mathbb{Z}\setminus A.$
A Chebyshev-type argument shows that at least $\tfrac{3}{4}$ of the mass of $g$ lies in $A$. Indeed, on $A^{\mathrm{c}}$ we have $(k-\mu)^{2}>R^{2}$, so
$J=\sum_{k}(k-\mu)^{2}g(k)\geq\sum_{k\in A^{\mathrm{c}}}(k-\mu)^{2}g(k)>R^{2}\sum_{k\in A^{\mathrm{c}}}g(k),$
hence
$\sum_{k\in A^{\mathrm{c}}}g(k)\leq\frac{J}{R^{2}}=\frac{J}{4J/G}=\frac{G}{4}.$
Therefore
$\sum_{k\in A}g(k)=G-\sum_{k\in A^{\mathrm{c}}}g(k)\geq\frac{3G}{4}.$
Next, we extract the “heavy” atom in $A$ via a cardinality bound and the pigeonhole principle:
$|A|\leq 2R+1=4\sqrt{\frac{J}{G}}+1,$
and so
$\sup_{k\in\mathbb{Z}}g(k)\ \geq\ \sup_{k\in A}g(k)\ \geq\ \frac{1}{|A|}\sum_{k\in A}g(k)\ \geq\ \frac{(3G/4)}{4\sqrt{J/G}+1}.$
The latter simplifies to our claimed bound. ∎
As above, $(\textup{{\bf p}},\textup{{\bf q}})$ is a dominating pair with $p_{i}\geq q_{i}$ and associated Poisson binomials $S_{\textup{{\bf p}}}$, $S_{\textup{{\bf q}}}$. We will apply the pigeonhole lemma to extract a “wedge” event with sufficient separation in probability under $S_{\textup{{\bf p}}}$ and $S_{\textup{{\bf q}}}$ to achieve the requisite TV lower bound. To this end, we define the $G$ and $J$ functionals — both in terms of the non-negative (by the monotone coupling) measure $g$ on $\mathbb{Z}$:
$g(k):=\mathbb{P}(X\geq k)-\mathbb{P}(Y\geq k),\qquad k\in\mathbb{Z}.$ (10)
Then $G=\sum_{k\in\mathbb{Z}}g(k)$ has the more familiar form
$G:=\Delta=m_{\textup{{\bf p}}}-m_{\textup{{\bf q}}}=\mathbb{E}X-\mathbb{E}Y,$
where $X:=S_{\textup{{\bf p}}},Y:=S_{\textup{{\bf q}}}$ and $m_{\textup{{\bf p}}}:=\mathbb{E}X,m_{\textup{{\bf q}}}:=\mathbb{E}Y$. Moreover, for any $k$,
$|g(k)|=|\mathbb{P}(X\geq k)-\mathbb{P}(Y\geq k)|\leq\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}}),$
and so
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\;\geq\;\sup_{k\in\mathbb{Z}}g(k).$ (11)
Finally, $J=\sum_{k\in\mathbb{Z}}(k-m_{\textup{{\bf p}}})^{2}g(k)$ is defined as in Lemma 7. The technical core of the argument hinges on upper bounding $J$:
Lemma 8.
Let $X,Y,\sigma_{\textup{{\bf p}}}^{2}=\operatorname{Var}(X),g,m_{\textup{{\bf p}}},\Delta,J$ be defined as above. Then
$J\;\leq\;2\,\Delta\,\bigl(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}\bigr).$
Proof.
For $i\in[n]$, write $\Delta_{i}:=p_{i}-q_{i}\geq 0$ with $\sum_{i}\Delta_{i}=\Delta$. For $t\in[0,1]$ set
$r_{i}(t):=q_{i}+t\Delta_{i},\quad X_{i}(t)\sim\operatorname{Ber}(r_{i}(t))\ \text{independent},\quad S(t):=\sum_{i=1}^{n}X_{i}(t),\quad T_{i}(t):=S(t)-X_{i}(t).$
Then $S(0)=Y$ and $S(1)=X$ in distribution. For fixed $n$, each $S(t)$ has a distribution that is a finite polynomial in the parameters $\{r_{i}(t)\}$, so for each $k$ and $i$ the function
$F_{k}(t)$ $:=$ $\mathbb{P}(S(t)\geq k)$
$=$ $r_{i}(t)\,\mathbb{P}(T_{i}(t)\geq k-1)\;+\;(1-r_{i}(t))\,\mathbb{P}(T_{i}(t)\geq k).$
is a polynomial in $t$ and hence differentiable on $[0,1]$. Differentiating,
$\frac{\partial}{\partial r_{i}}F_{k}=\mathbb{P}(T_{i}(t)\geq k-1)-\mathbb{P}(T_{i}(t)\geq k)=\mathbb{P}(T_{i}(t)=k-1).$
By the chain rule and the fact that $r_{i}^{\prime}(t)=\Delta_{i}$,
$\frac{\mathrm{d}}{\mathrm{d}t}\,\mathbb{P}(S(t)\geq k)=\sum_{i=1}^{n}\Delta_{i}\,\mathbb{P}\bigl(T_{i}(t)=k-1\bigr),\qquad k\in\mathbb{Z}.$
Integrating from 0 to $1$ and using $S(0)=Y$, $S(1)=X$ in distribution,
$g(k)=\mathbb{P}(X\geq k)-\mathbb{P}(Y\geq k)=\int_{0}^{1}\sum_{i=1}^{n}\Delta_{i}\,\mathbb{P}\bigl(T_{i}(t)=k-1\bigr)\,\mathrm{d}t,\qquad k\in\mathbb{Z}.$
In this formulation, the functional $J=\sum_{k}(k-m_{\textup{{\bf p}}})^{2}g(k)$ becomes
$J$ $=\int_{0}^{1}\sum_{i=1}^{n}\Delta_{i}\sum_{k\in\mathbb{Z}}(k-m_{\textup{{\bf p}}})^{2}\mathbb{P}\bigl(T_{i}(t)=k-1\bigr)\,\mathrm{d}t.$
Changing variable j=k1j=k-1,
$\sum_{k\in\mathbb{Z}}(k-m_{\textup{{\bf p}}})^{2}\mathbb{P}\bigl(T_{i}(t)=k-1\bigr)=\mathbb{E}_{t}\bigl[(T_{i}(t)+1-m_{\textup{{\bf p}}})^{2}\bigr],$
where $\mathbb{E}_{t}$ denotes expectation under the product law with parameters $\{r_{j}(t)\}_{j}$. Hence
$J=\int_{0}^{1}\sum_{i=1}^{n}\Delta_{i}\,\mathbb{E}_{t}\bigl[(T_{i}(t)+1-m_{\textup{{\bf p}}})^{2}\bigr]\,\mathrm{d}t.$ (12)
We bound $\mathbb{E}_{t}[(T_{i}(t)+1-m_{\textup{{\bf p}}})^{2}]$ uniformly in $t,i$ via the decomposition $\mathbb{E}[(Z-\alpha)]^{2}=\operatorname{Var}(Z)+(\mathbb{E}Z-\alpha)^{2}$, where $Z=T_{i}(t)+1$ and $\alpha=m_{\textup{{\bf p}}}$.
To bound the variance term, note that f(x)=x(1x)f(x)=x(1-x) is $1$-Lipschitz on $[0,1]$, and so
$|f(r_{j}(t))-f(p_{j})|\leq|r_{j}(t)-p_{j}|=(1-t)\Delta_{j}.$
Hence
$\operatorname{Var}_{t}(S(t))=\sum_{j}f(r_{j}(t))\leq\sum_{j}f(p_{j})+(1-t)\sum_{j}\Delta_{j}=\sigma_{\textup{{\bf p}}}^{2}+(1-t)\Delta\leq\sigma_{\textup{{\bf p}}}^{2}+\Delta\leq\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}.$
Since $T_{i}(t)=S(t)-X_{i}(t)$ with $X_{i}(t)$ Bernoulli,
$\operatorname{Var}_{t}(T_{i}(t))\leq\operatorname{Var}_{t}(S(t))\leq\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}.$ (13)
To bound the bias term, we have
$\mathbb{E}_{t}[S(t)]=m_{\textup{{\bf q}}}+t\Delta,$
and $m_{\textup{{\bf p}}}=m_{\textup{{\bf q}}}+\Delta$, so $\mathbb{E}_{t}[S(t)]-m_{\textup{{\bf p}}}=-(1-t)\Delta$. Also $r_{i}(t)=q_{i}+t\Delta_{i}$, so
$\mathbb{E}_{t}[T_{i}(t)]=m_{\textup{{\bf q}}}+t\Delta-(q_{i}+t\Delta_{i}),$
and hence
$M_{i}(t):=\mathbb{E}_{t}[T_{i}(t)+1-m_{\textup{{\bf p}}}]=1-q_{i}-\Delta+t(\Delta-\Delta_{i}).$
It is straightforward to see that $|M_{i}(t)|\leq\max(1,\Delta)$ and thus,
$\mathbb{E}_{t}\bigl[(T_{i}(t)+1-m_{\textup{{\bf p}}})^{2}\bigr]$ $=$ $\operatorname{Var}_{t}(T_{i}(t))+M_{i}(t)^{2}$
$\leq$ $(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2})+(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2})=2(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}).$
Substituting into (12) gives
$J\leq\int_{0}^{1}\sum_{i}\Delta_{i}\,2(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2})\,\mathrm{d}t=2\,\Delta\,(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}),$
as claimed. ∎
Corollary 9 (Proof of Theorem 2).
Let $(\textup{{\bf p}},\textup{{\bf q}})$ be a dominating pair with laws $S_{\textup{{\bf p}}},S_{\textup{{\bf q}}}$, sums $S_{\textup{{\bf p}}},S_{\textup{{\bf q}}}$, $\sigma_{\textup{{\bf p}}}^{2}:=\operatorname{Var}(S_{\textup{{\bf p}}})$, and $\Delta:=\mathbb{E}S_{\textup{{\bf p}}}-\mathbb{E}S_{\textup{{\bf q}}}$. Then
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\;\geq\;c\,\min\Bigl(1,\frac{\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}}\Bigr),$
for some universal constant $c\geq\frac{1}{12}$.
Proof.
Recall $g(k):=\mathbb{P}(S_{\textup{{\bf p}}}\geq k)-\mathbb{P}(S_{\textup{{\bf q}}}\geq k)\geq 0$ and
$\Delta=\sum_{k}g(k),\qquad J=\sum_{k}(k-m_{\textup{{\bf p}}})^{2}g(k),\qquad m_{\textup{{\bf p}}}=\mathbb{E}S_{\textup{{\bf p}}}.$
By (11), $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\geq\sup_{k\in\mathbb{Z}}g(k)$. By Lemma 8,
$J\;\leq\;2\,\Delta\,(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}).$
Apply Lemma 7 with $G=\Delta$ to obtain
$\sup_{k}g(k)\;\geq\;\frac{3\Delta}{16\sqrt{2(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2})}+4}.$
We compare $\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}}$ with $\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}$ and consider the two cases. If $\Delta\leq\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}$, then $\Delta^{2}\leq\sigma_{\textup{{\bf p}}}^{2}+1$, so
$\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}\leq 2(\sigma_{\textup{{\bf p}}}^{2}+1),$
and
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\geq\frac{3\Delta}{16\sqrt{2}\sqrt{2(\sigma_{\textup{{\bf p}}}^{2}+1)}+4}=\frac{3\Delta}{32\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+4}$
$\geq\frac{3\Delta}{36\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}}=\frac{\Delta}{12\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}}.$
If $\Delta\geq\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}$, then $\Delta^{2}\geq\sigma_{\textup{{\bf p}}}^{2}+1$, so
$\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}\leq 2\Delta^{2},$
and
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\geq\frac{3\Delta}{16\sqrt{2}\sqrt{2\Delta^{2}}+4}=\frac{3\Delta}{32\Delta+4}\geq\frac{3\Delta}{36\Delta}=\frac{1}{12}.$
Combining the two cases proves the claim. ∎
2.3 Proof of Theorem 3
Both inequalities invoke the data processing inequality; the first in the form
$\operatorname{TV}(P\otimes P^{\prime},Q\otimes Q^{\prime})$ $\geq$ $\max(\operatorname{TV}(P,Q),\operatorname{TV}(P^{\prime},Q^{\prime})),$
and the second in the form $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$. Given these, the claim is an immediate consequence of Theorem 2.
2.4 Proof of Lemma 4
Proof.
We begin by representing $\operatorname{Bin}(n,\bar{p}_{N})$ as a mixture. Let $M\sim\operatorname{Bin}(n,w)$ where $w=|I|/n$. Conditionally on $M=m$, let
$U_{m}\sim\operatorname{Bin}(m,\bar{p}_{I}),\qquad V_{m}\sim\operatorname{Bin}(n-m,\bar{p}_{J})$
be independent and set $S_{m}:=U_{m}+V_{m}$. We claim that $S_{M}\sim\operatorname{Bin}(n,\bar{p}_{N})$. Indeed, for $t\in\mathbb{R}$, put $s:=1-\bar{p}_{I}+\bar{p}_{I}t$ and $r:=1-\bar{p}_{J}+\bar{p}_{J}t$. Compute the probability generating function conditional on $M$:
$\mathbb{E}[t^{S_{M}}\mid M]=\mathbb{E}[t^{U_{M}+V_{M}}\mid M]=\mathbb{E}[t^{U_{M}}\mid M]\;\mathbb{E}[t^{V_{M}}\mid M]=s^{\,M}r^{\,n-M},$
so $\mathbb{E}[t^{S_{M}}]=\mathbb{E}[s^{M}r^{n-M}]$. Represent $M=\sum_{i=1}^{n}X_{i}$ where $X_{1},\dots,X_{n}$ are i.i.d. $\operatorname{Ber}(w)$. Then $n-M=\sum_{i=1}^{n}(1-X_{i})$ and
$s^{M}r^{n-M}=\prod_{i=1}^{n}s^{X_{i}}r^{1-X_{i}}=\prod_{i=1}^{n}\bigl(r+(s-r)X_{i}\bigr).$
Taking expectation and using independence,
$\mathbb{E}[s^{M}r^{n-M}]=\prod_{i=1}^{n}\mathbb{E}\bigl[r+(s-r)X_{i}\bigr]=\bigl(r+(s-r)\mathbb{E}[X_{1}]\bigr)^{n}=\bigl((1-w)r+ws\bigr)^{n}.$
Substituting $r=1-\bar{p}_{J}+\bar{p}_{J}t$ and $s=1-\bar{p}_{I}+\bar{p}_{I}t$ gives
$\mathbb{E}[t^{S_{M}}]=\bigl(w(1-\bar{p}_{I}+\bar{p}_{I}t)+(1-w)(1-\bar{p}_{J}+\bar{p}_{J}t)\bigr)^{n}=(1-\bar{p}_{N}+\bar{p}_{N}\,t)^{n},$
which is the PGF of $\operatorname{Bin}(n,\bar{p}_{N})$. Since a distribution with finite support is characterized by its PGF, $S_{M}\sim\operatorname{Bin}(n,\bar{p}_{N})$. Similarly, if $U^{\prime}_{m}\sim\operatorname{Bin}(m,\bar{q}_{I})$ and $V^{\prime}_{m}\sim\operatorname{Bin}(n-m,\bar{q}_{J})$ are independent and $S^{\prime}_{m}:=U^{\prime}_{m}+V^{\prime}_{m}$, then the same calculation yields $S^{\prime}_{M}\sim\operatorname{Bin}(n,\bar{q}_{N})$. Therefore
$\delta_{N}=\operatorname{TV}(S_{M},S^{\prime}_{M}).$
For any event $A\subset\{0,1,\dots,n\}$,
$\mathbb{P}(S_{M}\in A)-\mathbb{P}(S^{\prime}_{M}\in A)=\mathbb{E}\!\left[\mathbb{P}(S_{M}\in A\mid M)-\mathbb{P}(S^{\prime}_{M}\in A\mid M)\right]\leq\mathbb{E}\!\left[\operatorname{TV}(\operatorname{\mathcal{L}}(S_{M}\mid M),\operatorname{\mathcal{L}}(S^{\prime}_{M}\mid M))\right],$
and taking the supremum over $A$ yields
$\operatorname{TV}(S_{M},S^{\prime}_{M})\leq\mathbb{E}\!\left[\operatorname{TV}(\operatorname{\mathcal{L}}(S_{M}\mid M),\operatorname{\mathcal{L}}(S^{\prime}_{M}\mid M))\right]=\sum_{m=0}^{n}\mathbb{P}(M=m)\,\operatorname{TV}(S_{m},S^{\prime}_{m}).$
Next, the map $(u,v)\mapsto u+v$ is a Markov kernel, and so
$\operatorname{TV}(S_{m},S^{\prime}_{m})\leq\operatorname{TV}\bigl((U_{m},V_{m}),(U^{\prime}_{m},V^{\prime}_{m})\bigr).$
Moreover, for product measures one has
$\operatorname{TV}(P\times R,\;Q\times S)\leq\operatorname{TV}(P,Q)+\operatorname{TV}(R,S),$ (14)
whence
$\operatorname{TV}(S_{m},S^{\prime}_{m})\leq\operatorname{TV}(U_{m},U^{\prime}_{m})+\operatorname{TV}(V_{m},V^{\prime}_{m}).$
Define
$\Delta_{I}(m):=\operatorname{TV}(\operatorname{Bin}(m,\bar{p}_{I}),\operatorname{Bin}(m,\bar{q}_{I})),\qquad\Delta_{J}(k):=\operatorname{TV}(\operatorname{Bin}(k,\bar{p}_{J}),\operatorname{Bin}(k,\bar{q}_{J})).$
Then
$\delta_{N}\leq\mathbb{E}[\Delta_{I}(M)]+\mathbb{E}[\Delta_{J}(n-M)].$
It remains to estimate the two expectations. Fix $\theta,\theta^{\prime}\in[0,1]$ and define $f:\mathbb{N}\to\mathbb{R}$ by
$f(m):=\operatorname{TV}(\operatorname{Bin}(m,\theta),\operatorname{Bin}(m,\theta^{\prime})).$
We claim that $f$ is nondecreasing and subadditive. To show subadditivity, let $X\sim\operatorname{Bin}(m,\theta)$, $Y\sim\operatorname{Bin}(k,\theta)$ be independent so that $X+Y\sim\operatorname{Bin}(m+k,\theta)$, and similarly let $X^{\prime}\sim\operatorname{Bin}(m,\theta^{\prime})$, $Y^{\prime}\sim\operatorname{Bin}(k,\theta^{\prime})$ independent so that $X^{\prime}+Y^{\prime}\sim\operatorname{Bin}(m+k,\theta^{\prime})$. Then
$f(m+k)=\operatorname{TV}(X+Y,\;X^{\prime}+Y^{\prime})\leq\operatorname{TV}(X,X^{\prime})+\operatorname{TV}(Y,Y^{\prime})=f(m)+f(k).$
To show monotonicity, observe that the following is a Markov kernel that maps $\operatorname{Bin}(m+1,\theta)$ to $\operatorname{Bin}(m,\theta)$: interpret $\operatorname{Bin}(m+1,\theta)$ as the number of successes in $m+1$ i.i.d. $\operatorname{Ber}(\theta)$ trials and delete one uniformly random trial. Conditional on seeing $k$ successes among $m+1$ trials, the remaining number of successes equals $k$ with probability $\frac{m+1-k}{m+1}$ (deleted a failure) and equals k1k-1 with probability $\frac{k}{m+1}$ (deleted a success). This kernel does not depend on $\theta$, and it sends $\operatorname{Bin}(m+1,\theta)$ to $\operatorname{Bin}(m,\theta)$ for every $\theta$. Therefore, by TV contraction under the same kernel,
$f(m)=\operatorname{TV}(T_{m}(\operatorname{Bin}(m+1,\theta)),T_{m}(\operatorname{Bin}(m+1,\theta^{\prime})))\leq\operatorname{TV}(\operatorname{Bin}(m+1,\theta),\operatorname{Bin}(m+1,\theta^{\prime}))=f(m+1).$
Now let $m_{0}\geq 1$ be an integer. By subadditivity and monotonicity, for every $m\geq 0$, writing $m=km_{0}+r$ with $0\leq r<m_{0}$,
$f(m)\leq kf(m_{0})+f(r)\leq(k+1)f(m_{0})=\left\lceil\frac{m}{m_{0}}\right\rceil f(m_{0}).$
Hence, for any integer–valued $M$ with $\mathbb{E}M=m_{0}$, using $\lceil x\rceil\leq x+1$,
$\mathbb{E}[f(M)]\leq f(m_{0})\,\mathbb{E}\!\left[\left\lceil\frac{M}{m_{0}}\right\rceil\right]\leq f(m_{0})\,\mathbb{E}\!\left[\frac{M}{m_{0}}+1\right]=2f(m_{0}).$
Apply this with $f=\Delta_{I}$ and $m_{0}=|I|$ (note $\mathbb{E}M=nw=|I|$) to get
$\mathbb{E}[\Delta_{I}(M)]\leq 2\Delta_{I}(|I|)=2\delta_{I},$
and with $f=\Delta_{J}$ and $m_{0}=|J|$ (since $\mathbb{E}(n-M)=|J|$) to get
$\mathbb{E}[\Delta_{J}(n-M)]\leq 2\Delta_{J}(|J|)=2\delta_{J}.$
This completes the proof. ∎
Acknowledgments.
Some of the proofs were assisted by ChatGPT Pro; we thank Sébastien Bubeck for making this version available to us. We also thank Mark Sellke and Nikita Zhivotovskiy for pointing out errors in earlier versions.

### v2

@@
 Equivalently, if $X_{i}\sim\operatorname{Ber}(p_{i})$ are independent, then $S_{\textup{{\bf p}}}:=\sum_{i=1}^{n}X_{i}$ is distributed according to this law.
-The sum map is easily seen to be a Markov kernel, and hence the Data Processing Inequality [12, Theorem 7.4] applies: $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$, where $\operatorname{TV}$ is the total variation distance.
+The sum map is easily seen to be a Markov kernel, and hence the Data Processing Inequality [16, Theorem 7.4] applies: $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$, where $\operatorname{TV}$ is the total variation distance.
 Intuitively, since information is lost under the sum map, the distributions can only become closer to each other.
@@
 Since this operation also erases information, we might likewise expect it to decrease the TV distance.
-One immediate glitch is that the homogenization map cannot be realized by any Markov kernel.
+One immediate obstruction is that the homogenization map cannot be realized by any Markov kernel.
 Indeed, suppose to the contrary that for $n\geq 2$, some Markov kernel $K$ maps $X\sim\operatorname{Ber}(\textup{{\bf p}})$ to $Y\sim\operatorname{Ber}(\bar{p})^{\otimes n}$.
@@
 Then the law of $X$ is affine in $t$ (it is a $t$-convex combination of two atoms), as is the pushforward measure under $\tilde{K}$.
-However, $\mathbb{P}(Z=2)=\binom{n}{2}\left(\frac{t}{n}\right)^{2}\left(1-\frac{t}{n}\right)^{n-2}$, which is not affine in $t$.
+However, $\mathbb{P}(Z=2)=\binom{n}{2}\left(\frac{t}{n}\right)^{2}\left(1-\frac{t}{n}\right)^{n-2}$, which is not affine in tt.
+[^2] For the special case of $\textup{{\bf p}},\textup{{\bf q}}$ that are symmetric about $\frac{1}{2}$, meaning that $\textup{{\bf q}}=\textup{{\bf 1}}-\textup{{\bf p}}$, it turns out that the homogenization operator can be realized as a Markov kernel; see Appendix.
 This indicates that the standard information-theoretic argument for proving data processing inequalities does not apply, necessitating novel tools.
@@
 For general $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, only the upper bound $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\lesssim\Phi(\textup{{\bf p}},\textup{{\bf q}})$ holds, which can be sharpened to $\min(\Phi(\textup{{\bf p}},\textup{{\bf q}}),\Phi(\textup{{\bf q}},\textup{{\bf p}}))$ by symmetry.
-Unlike $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$, whose exact computation is infeasible (see below), $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ is computable exactly in time $O(n^{2})$ via a simple recurrence [15, Section 5].
+Unlike $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$, whose exact computation is infeasible (see below), $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ is computable exactly in time $O(n^{2})$ via a simple recurrence [20, Section 5].
 Still, one should not expect any analytically tractable analogue of (2) for general $\textup{{\bf p}},\textup{{\bf q}}$ due to cancellations and convolutional smoothing.
@@
 Numerically (see [1, Remark 1]), $\eta_{\mathrm{BCV}}\approx 0.4688223555$ and hence $C_{\mathrm{BCV}}\approx 1.2123507747$.
-Theorem 1.
+Theorem 1.1.
 For $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ and $\Phi$ as in (1),
-$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\leq$ $C_{\mathrm{BCV}}\,\min(\Phi(\textup{{\bf p}},\textup{{\bf q}}),\Phi(\textup{{\bf q}},\textup{{\bf p}})).$
+$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\leq$ $2C_{\mathrm{BCV}}\,\min(\Phi(\textup{{\bf p}},\textup{{\bf q}}),\Phi(\textup{{\bf q}},\textup{{\bf p}})).$
 The technical core of the paper is a matching lower bound for dominating pairs:
-Theorem 2.
-For $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ satisfy $p_{i}\geq q_{i}$, $i\in[n]$, and $\Phi$ as in (1),
-$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\geq$ $\frac{1}{12}\Phi(\textup{{\bf p}},\textup{{\bf q}}).$
-The assumption $\textup{{\bf p}}\geq\textup{{\bf q}}$ is essential for the lower bound, as illustrated by the example $\textup{{\bf p}}=(1,\,0,\,\tfrac{1}{2})$ and $\textup{{\bf q}}=(0,\,1,\,\tfrac{1}{2}+\varepsilon)$; here, $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})=\varepsilon$ while $\Phi(\textup{{\bf p}},\textup{{\bf q}})=1$.
+Theorem 1.2.
+For $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ satisfying $p_{i}\geq q_{i}$, $i\in[n]$, and $\Phi$ as in (1),
+$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\geq$ $\frac{1}{9}\Phi(\textup{{\bf p}},\textup{{\bf q}}).$
+The assumption $\textup{{\bf p}}\geq\textup{{\bf q}}$ is essential for the lower bound, as illustrated by the example $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, for even $n$, given by
+$\textup{{\bf p}}=\Bigl(\tfrac{1}{2},\tfrac{1}{2},\dots,\tfrac{1}{2}\Bigr)$, $\textup{{\bf q}}=\Bigl(\tfrac{1}{2}+\delta,\dots,\tfrac{1}{2}+\delta,\ \tfrac{1}{2}-\delta,\dots,\tfrac{1}{2}-\delta\Bigr),$
+with $n/2$ copies of $\tfrac{1}{2}+\delta$ followed by $n/2$ copies of $\tfrac{1}{2}-\delta$.
+The choice
+$\delta=1/\sqrt{n}$ yields $\Phi(\textup{{\bf p}},\textup{{\bf q}})=1$, but $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\asymp 1/n$.
 In general, for any $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ and any partition $I,J$ of $[n]$, the trivial estimate via the partitioned pairs $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\leq\operatorname{TV}(S_{\textup{{\bf p}}_{I}},S_{\textup{{\bf q}}_{I}})+\operatorname{TV}(S_{\textup{{\bf p}}_{J}},S_{\textup{{\bf q}}_{J}})$ holds.
 However, this method cannot yield a matching lower bound, as illustrated by the example $a,b=\tfrac{1}{2}\pm\varepsilon$, $\textup{{\bf p}}=(a,a,b,b)$, $\textup{{\bf q}}=(a,b,a,b)$ with the partition $I=\left\{1,2\right\}$, $J=\left\{3,4\right\}$.
-Although for general $\textup{{\bf p}},\textup{{\bf q}}$, the approach of Theorem 2 does not lower-bound $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$, it does yield a lower bound for the larger quantity $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$:
-Theorem 3.
+Although for general $\textup{{\bf p}},\textup{{\bf q}}$, the approach of Theorem 1.2 does not lower-bound $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$, it does yield a lower bound for the larger quantity $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$:
+Theorem 1.3.
 For $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, let $I=\left\{i\in[n]:p_{i}\geq q_{i}\right\}$ and $J=[n]\setminus I$, with $\textup{{\bf p}}_{I},\textup{{\bf p}}_{J},\textup{{\bf q}}_{I},\textup{{\bf q}}_{J}$ the corresponding subsequences of $\textup{{\bf p}},\textup{{\bf q}}$.
@@
 $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\geq$ $\max\left(\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}_{I}),\operatorname{Ber}(\textup{{\bf q}}_{I})),\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}_{J}),\operatorname{Ber}(\textup{{\bf q}}_{J}))\right)$
-$\geq$ $\frac{1}{12}\max(\Phi(\textup{{\bf p}}_{I},\textup{{\bf q}}_{I}),\Phi(\textup{{\bf q}}_{J},\textup{{\bf p}}_{J})).$
+$\geq$ $\frac{1}{9}\max(\Phi(\textup{{\bf p}}_{I},\textup{{\bf q}}_{I}),\Phi(\textup{{\bf q}}_{J},\textup{{\bf p}}_{J})).$
 The last piece to complete our program is a homogenization result for binomials:
-Lemma 4.
+Lemma 1.4.
 For $n\geq 1$, $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, and $\emptyset\neq A\subset N:=[n]$, define $\bar{p}_{A}=|A|^{-1}\sum_{i\in A}p_{i}$ and $\bar{q}_{A}$ analogously; put
@@
 The homogenization inequality claimed in (3) then follows:
-Theorem 5.
+Theorem 1.5.
 For all $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, we have
 $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\geq$ $c\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q})),$
-where $c\geq\frac{1}{48C_{\mathrm{BCV}}}\approx 0.017$ is a universal constant.
+where $c\geq\frac{1}{72C_{\mathrm{BCV}}}\approx 0.0115$ is a universal constant.
 Remark.
+As the remark preceding Theorem 2.1 indicates, a sharpened constant $c\geq\frac{1}{36C_{\mathrm{BCV}}}$ is straightforward to obtain via analogous arguments.
 The choice $\textup{{\bf p}}=(1-2\varepsilon,\tfrac{1}{2})$, $\textup{{\bf q}}=(1,\tfrac{1}{2}+\varepsilon)$, $\varepsilon\to 0$ demonstrates that it is possible for homogenization to (slightly) increase the TV distance, and also that $8/9$ is an upper bound on $c$.
@@
 Proof.
-Partition $[n]$ into $I,J$ as in Theorem 3 and ignore the trivial case where one of $I,J$ is empty. Then Theorem 3 implies
-$\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\geq$ $\frac{1}{12}\max(\Phi(\textup{{\bf p}}_{I},\textup{{\bf q}}_{I}),\Phi(\textup{{\bf q}}_{J},\textup{{\bf p}}_{J}))$
-$\geq$ $\frac{1}{24}\left(\Phi(\textup{{\bf p}}_{I},\textup{{\bf q}}_{I})+\Phi(\textup{{\bf q}}_{J},\textup{{\bf p}}_{J})\right).$
+Partition $[n]$ into $I,J$ as in Theorem 1.3 and ignore the trivial case where one of $I,J$ is empty. Then Theorem 1.3 implies
+$\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\geq$ $\frac{1}{9}\max(\Phi(\textup{{\bf p}}_{I},\textup{{\bf q}}_{I}),\Phi(\textup{{\bf q}}_{J},\textup{{\bf p}}_{J}))$
+$\geq$ $\frac{1}{18}\left(\Phi(\textup{{\bf p}}_{I},\textup{{\bf q}}_{I})+\Phi(\textup{{\bf q}}_{J},\textup{{\bf p}}_{J})\right).$
 A simple convexity argument shows that homogenization cannot increase $\Phi$:
@@
 where $\bar{p}\boldsymbol{1}=(\bar{p},\bar{p},\ldots,\bar{p})$ is of the same dimension as p.
-Now homogenize $\bar{p}_{I}$, $\bar{p}_{J}$, $\bar{q}_{I}$, $\bar{q}_{J}$ as in Lemma 4. Since $\operatorname{Bin}(|I|,\bar{p}_{I})$ is a special case of Poisson binomial, Theorem 1 applies:
+Now homogenize $\bar{p}_{I}$, $\bar{p}_{J}$, $\bar{q}_{I}$, $\bar{q}_{J}$ as in Lemma 1.4.
+Since $\operatorname{Bin}(|I|,\bar{p}_{I})$ is a special case of Poisson binomial, Theorem 1.1 applies:
 $\Phi(\textup{{\bf p}}_{I},\textup{{\bf q}}_{I})+\Phi(\textup{{\bf q}}_{J},\textup{{\bf p}}_{J})$ $\geq$ $\Phi(\bar{p}_{I}\boldsymbol{1},\bar{q}_{I}\boldsymbol{1})+\Phi(\bar{q}_{J}\boldsymbol{1},\bar{p}_{J}\boldsymbol{1})$
-$\geq$ $\frac{1}{C_{\mathrm{BCV}}}\left(\operatorname{TV}(\operatorname{Bin}(|I|,\bar{p}_{I}),\operatorname{Bin}(|I|,\bar{q}_{I}))+\operatorname{TV}(\operatorname{Bin}(|J|,\bar{p}_{J}),\operatorname{Bin}(|J|,\bar{q}_{J}))\right)$
-$\geq$ $\frac{1}{2C_{\mathrm{BCV}}}\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q})),$
-where the last inequality is by Lemma 4.
+$\geq$ $\frac{1}{2C_{\mathrm{BCV}}}\left(\operatorname{TV}(\operatorname{Bin}(|I|,\bar{p}_{I}),\operatorname{Bin}(|I|,\bar{q}_{I}))+\operatorname{TV}(\operatorname{Bin}(|J|,\bar{p}_{J}),\operatorname{Bin}(|J|,\bar{q}_{J}))\right)$
+$\geq$ $\frac{1}{4C_{\mathrm{BCV}}}\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q})),$
+where the last inequality is by Lemma 1.4.
 Finally, the identity $\operatorname{TV}(\operatorname{Ber}(\bar{p})^{\otimes n},\operatorname{Ber}(\bar{q})^{\otimes n})$ == $\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q}))$ is a standard consequence of the Neyman–Pearson lemma (the likelihood ratio is determined by and monotone in the sum).
@@
 Although the homogenization inequality relies on novel structural insights into the Poisson binomial, there are compelling reasons to believe that a great deal more structure remains to be uncovered, currently out of reach.
-Indeed, copious numerical trials unambiguously indicate that the correct constant in Theorem 5 should be $c=8/9$ and also that the bound in Lemma 4 can be sharpened to $\delta_{N}\leq\delta_{I}+\delta_{J}-\delta_{I}\delta_{J}$ (the latter, in particular, appears deceptively simple, since only binomials are involved).
+Indeed, extensive numerical experiments suggest that the correct constant in Theorem 1.5 should be $c=8/9$ and also that the bound in Lemma 1.4 can be sharpened to $\delta_{N}\leq\delta_{I}+\delta_{J}-\delta_{I}\delta_{J}$ (the latter, in particular, appears deceptively simple, since only binomials are involved).
 Our present methods do not seem to provide any pathway to these conjecturally optimal bounds, which will have to await further structural advances.
 Related work.
-This paper follows a program initiated by [8] to provide simple, analytically tractable upper and lower estimates on $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ in terms of the $\textup{{\bf p}},\textup{{\bf q}}$; the latter’s contribution was a sharpening of the trivial lower bound $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{\infty}$ to $\gtrsim\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{2}$.
-We note that generally, the latter is incomparable with the lower bound $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\gtrsim\Phi(\textup{{\bf p}},\textup{{\bf q}})$ implied by Theorem 2.
+This paper follows a program initiated by [9] to provide simple, analytically tractable upper and lower estimates on $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ in terms of the $\textup{{\bf p}},\textup{{\bf q}}$; the latter’s contribution consisted of sharpening of the trivial lower bound $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{\infty}$ to $\gtrsim\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{2}$.
+We note that generally, this result is incomparable with the lower bound $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\gtrsim$ $\Phi(\textup{{\bf p}},\textup{{\bf q}})$ implied by Theorem 1.2 for $\textup{{\bf p}}\geq\textup{{\bf q}}$.
 Indeed, the choice $q_{i}\equiv\frac{1}{2}$ for $i\in[n]$ and $p_{i}=1/2$, $i>1$ and $p_{1}=\frac{1}{2}+\varepsilon$ yields $\Phi(\textup{{\bf p}},\textup{{\bf q}})\asymp\varepsilon/\sqrt{n}$ while $\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{2}=\varepsilon$.
 For the other direction, the choice $\textup{{\bf q}}\equiv 0$ and $\textup{{\bf p}}\equiv 1/n$ yields $\Phi(\textup{{\bf p}},\textup{{\bf q}})\asymp 1$ and $\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|=n^{-1/2}$.
+This program is continued in [10], where an analytical closed-form $O(\sqrt{\log n})$-factor approximation to $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ is obtained.
 Homogenization inequalities appear not to have been widely studied; one classic result due to Hoeffding [8] is that homogenization under a fixed-mean constraint maximizes $\mathbb{E}g(S_{\textup{{\bf p}}})$ for any convex $g$.
-The study of $S_{\textup{{\bf p}}}$ is amply motivated by numerous applications and a venerable line of work has considered approximating this distribution by simpler simpler ones [15]; perhaps most famous is Le Cam’s inequality $\operatorname{TV}(S_{\textup{{\bf p}}},\mathrm{Poi}(\sum p_{i}))\leq\sum{p_{i}^{2}}$ [9].
+The study of $S_{\textup{{\bf p}}}$ is amply motivated by numerous applications and a venerable line of work has considered approximating this distribution by simpler ones [20]; perhaps most famous is Le Cam’s inequality $\operatorname{TV}(S_{\textup{{\bf p}}},\mathrm{Poi}(\sum p_{i}))\leq\sum{p_{i}^{2}}$ [11].
 Note, however, that the latter yields an additive, rather than multiplicative approximation to $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$.
-Binomial approximation bounds in strong metrics (including $\operatorname{TV}$) appear in [4] and refinements based on orthogonal-polynomial expansions and asymptotics in [13].
+Binomial approximation bounds in strong metrics (including $\operatorname{TV}$) appear in [5] and refinements based on orthogonal-polynomial expansions and asymptotics in [17], then generalized to arbitrary laws [18] and also gave an additive homogenization-type estimate.
 Later, [19] obtained TV bounds for approximating general convolutions by compound Poisson laws using Kerstan’s method together with novel “smoothness inequalities”.
 A delicate structural result due to [4] (also ultimately yielding an additive estimate) shows that every Poisson binomial admits an $\varepsilon$-approximation in TV by a distribution of one of two canonical “compressed” types: either a sparse $S_{\textup{{\bf p}}}$ with only $O(1/\varepsilon^{3})$ nontrivial summands, or a near-binomial form.
-Of particular relevance to this work is [1], whose structural results we build upon in Theorem 1.
+Of particular relevance to this work is [1], whose structural results we build upon in Theorem 1.1.
 More generally, $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ admits classic approximations in terms of more analytically tractable proxies such as KL-divergence and the Hellinger distance; their properties and limitations are discussed in [9].
@@
 2 Proofs
-2.1 Proof of Theorem 1
-We will prove a sharper, symmetric bound, which immediately implies the one in Theorem 1:
-Theorem 6.
+2.1 Proof of Theorem 1.1
+We will prove a more general, symmetric bound, which immediately implies the one in Theorem 1.1.
+The constant $2C_{\mathrm{BCV}}$ in the latter’s bound could be sharpened by a factor of $2$ via a direct (asymmetric) proof.
+L.
+Mattner [12] has brought to our attention the result of [21], which upper bounds $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ by
+$\frac{5|\mathbb{E}S_{\textup{{\bf p}}}-\mathbb{E}S_{\textup{{\bf q}}}|}{\sqrt{\sigma^{2}_{\textup{{\bf p}}}+\sigma^{2}_{\textup{{\bf q}}}}}+\frac{12|\mathbb{E}S_{\textup{{\bf p}}}-\mathbb{E}S_{\textup{{\bf q}}}|}{{1+\sigma^{2}_{\textup{{\bf p}}}+\sigma^{2}_{\textup{{\bf q}}}}}.$
+This estimate has the attractive property of being topologically equivalent to $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ in the sense that the two converge to 0 on precisely the same set of sequence pairs.
+It is currently an open question whether a version of our $\Phi$ functional could be made topologically equivalent to $\operatorname{TV}$ as well.
+Theorem 2.1.
 For $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ and $S_{\textup{{\bf p}}},S_{\textup{{\bf q}}}$ the corresponding Poisson binomials with variances $\sigma_{\textup{{\bf p}}}^{2}$, $\sigma_{\textup{{\bf q}}}^{2}$, we have
@@
 by a simple telescoping argument.
-Anti-concentration for Poisson binomials (Baillon–Cominetti–Vaisman). If $Z=\sum_{j=1}^{m}\operatorname{Ber}(r_{j})$ has variance $v=\operatorname{Var}(Z)$, then [1, Theorem 1] proves the sharp bound
+[^3] M.
+Lutz [12] informs us that this result is contained in [2] with a slightly worse constant.
+Anti-concentration for Poisson binomials (Baillon–Cominetti–Vaisman).
+If $Z=\sum_{j=1}^{m}\operatorname{Ber}(r_{j})$ has variance $v=\operatorname{Var}(Z)$, then [1, Theorem 1] proves the sharp bound
 $\max_{k}\mathbb{P}(Z=k)\leq\frac{\eta_{\mathrm{BCV}}}{\sqrt{v}}\qquad(v>0),$ (6)
@@
 $\max_{k}\mathbb{P}(Z=k)\leq\min\!\left\{1,\ \frac{\eta_{\mathrm{BCV}}}{\sqrt{v}}\right\}.$ (7)
-Unimodality of Poisson binomial pmfs. It is a classic consequence of the Aissen-Endrei-Schoenberg-Whitney theorem [15, Section 4] that the Poisson binomial has a log-concave, and hence unimodal, law.
-To combine these three ingredients, we begin with the interpolation. For $t\in[0,1]$ define
+Unimodality of Poisson binomial pmfs. It is a classic consequence of the Aissen-Edrei-Schoenberg-Whitney theorem [20, Section 4] that the Poisson binomial
+270]
+has a log-concave, and hence unimodal, law. To combine these three ingredients, we begin with the interpolation. For $t\in[0,1]$ define
 $r_{i}(t):=(1-t)q_{i}+tp_{i},\qquad S(t):=\sum_{i=1}^{n}\operatorname{Ber}(r_{i}(t)),\qquad t\in[0,1],$
@@
 $T_{i}(t):=S(t)-X_{i}(t)=\sum_{j\neq i}\operatorname{Ber}(r_{j}(t)).$
-Then $T_{i}(t)$ is independent of $X_{i}(t)$ and $S(t)=T_{i}(t)+X_{i}(t)$.
-Conditioning on $X_{i}(t)$ gives
+Then $T_{i}(t)$ is independent of $X_{i}(t)$ and $S(t)=T_{i}(t)+X_{i}(t)$. Conditioning on $X_{i}(t)$ gives
 $f_{A}(t)=r_{i}(t)\mathbb{P}(T_{i}(t)+1\in A)+(1-r_{i}(t))\mathbb{P}(T_{i}(t)\in A).$
@@
 $\bigl|\mathbb{P}(S_{\textup{{\bf p}}}\in A)-\mathbb{P}(S_{\textup{{\bf q}}}\in A)\bigr|\leq\frac{2C_{\mathrm{BCV}}\,\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+\sqrt{\sigma_{\textup{{\bf q}}}^{2}+1}}.$
-Taking $\sup_{A}$ proves (6). ∎
-2.2 Proof of Theorem 2
-We begin with an abstract result showing that if a positive measure $g$ on $\mathbb{Z}$ has total mass $G$ and its variance-like spread about $\mu$ is controlled by $J<\infty$, this implies a lower bound on the mass of the heaviest atom.
-Lemma 7 (Pigeonhole).
+Taking $\sup_{A}$ proves (2.1). ∎
+2.2 Proof of Theorem 1.2
+We make use of a concentration-variance inequality [13, Disp.(18)], which immediately implies the following “pigeonhole” lemma: [^4] We thank M.
+Lutz [12] for bringing this to our attention and improving our original lower bound by a factor of $4/3$.
+Lemma 2.2.
 For $g:\mathbb{Z}\to[0,\infty)$ and $\mu\in\mathbb{R}$, define
@@
 If $0<G,J<\infty$, then
-$\sup_{k\in\mathbb{Z}}g(k)\;\geq\;\frac{3\,G^{3/2}}{16\sqrt{J}+4\sqrt{G}}.$
-Proof.
-Put $R:=2\sqrt{{J}/{G}}$ and define
-$A:=\{k\in\mathbb{Z}:\ |k-\mu|\leq R\},\qquad A^{\mathrm{c}}:=\mathbb{Z}\setminus A.$
-A Chebyshev-type argument shows that at least $\tfrac{3}{4}$ of the mass of $g$ lies in $A$. Indeed, on $A^{\mathrm{c}}$ we have $(k-\mu)^{2}>R^{2}$, so
-$J=\sum_{k}(k-\mu)^{2}g(k)\geq\sum_{k\in A^{\mathrm{c}}}(k-\mu)^{2}g(k)>R^{2}\sum_{k\in A^{\mathrm{c}}}g(k),$
-hence
-$\sum_{k\in A^{\mathrm{c}}}g(k)\leq\frac{J}{R^{2}}=\frac{J}{4J/G}=\frac{G}{4}.$
-Therefore
-$\sum_{k\in A}g(k)=G-\sum_{k\in A^{\mathrm{c}}}g(k)\geq\frac{3G}{4}.$
-Next, we extract the “heavy” atom in $A$ via a cardinality bound and the pigeonhole principle:
-$|A|\leq 2R+1=4\sqrt{\frac{J}{G}}+1,$
-and so
-$\sup_{k\in\mathbb{Z}}g(k)\ \geq\ \sup_{k\in A}g(k)\ \geq\ \frac{1}{|A|}\sum_{k\in A}g(k)\ \geq\ \frac{(3G/4)}{4\sqrt{J/G}+1}.$
-The latter simplifies to our claimed bound. ∎
+$\sup_{k\in\mathbb{Z}}g(k)\;\geq\;\frac{G^{3/2}}{4\sqrt{J}+\sqrt{G}}.$
+Proof (sketch).
+Normalize $g$ to be a probability measure $P$ via $P(\left\{k\right\})=g(k)/G$ and let $X$ be distributed according to $P$. Using the variational characterization of variance, $Var(X)\leq J/G$.
+By Mattner–Schulz [13, Disp. (18)], for any law with variance $\sigma^{2}$ and any $h>0$,
+$\sup_{x\in\mathbb{R}}P\bigl((x,x+h)\bigr)\;\geq\;\frac{h}{\sqrt{h^{2}+12\sigma^{2}}}.$
+Apply this with $h=1$ and $\sigma^{2}=\operatorname{Var}(X)$. Since $P$ is supported on $\mathbb{Z}$, any interval of length $1$ contains at most one atom, and hence
+$\sup_{x\in\mathbb{R}}P((x,x+1))=\max_{k\in\mathbb{Z}}P(\{k\})=\frac{1}{G}\max_{k\in\mathbb{Z}}g(k).$
+Therefore,
+$\frac{1}{G}\max_{k}g(k)\;\geq\;\frac{1}{\sqrt{1+12\operatorname{Var}(X)}}.$
+Using $\operatorname{Var}(X)\leq J/G$,
+$\max_{k}g(k)\;\geq\;\frac{G}{\sqrt{1+12(J/G)}}=\frac{G^{3/2}}{\sqrt{G+12J}}.$
+Finally, since $\sqrt{G+12J}\leq\sqrt{G}+4\sqrt{J}$, we conclude
+$\max_{k\in\mathbb{Z}}g(k)\;\geq\;\frac{G^{3/2}}{\,4\sqrt{J}+\sqrt{G}\,}$,
+which is the claimed bound. ∎
 As above, $(\textup{{\bf p}},\textup{{\bf q}})$ is a dominating pair with $p_{i}\geq q_{i}$ and associated Poisson binomials $S_{\textup{{\bf p}}}$, $S_{\textup{{\bf q}}}$.
@@
 Then $G=\sum_{k\in\mathbb{Z}}g(k)$ has the more familiar form
-$G:=\Delta=m_{\textup{{\bf p}}}-m_{\textup{{\bf q}}}=\mathbb{E}X-\mathbb{E}Y,$
-where $X:=S_{\textup{{\bf p}}},Y:=S_{\textup{{\bf q}}}$ and $m_{\textup{{\bf p}}}:=\mathbb{E}X,m_{\textup{{\bf q}}}:=\mathbb{E}Y$. Moreover, for any $k$,
+$G:=\Delta=\mu_{\textup{{\bf p}}}-\mu_{\textup{{\bf q}}}=\mathbb{E}X-\mathbb{E}Y,$
+where $X:=S_{\textup{{\bf p}}},Y:=S_{\textup{{\bf q}}}$ and $\mu_{\textup{{\bf p}}}:=\mathbb{E}X,\mu_{\textup{{\bf q}}}:=\mathbb{E}Y$. Moreover, for any $k$,
 $|g(k)|=|\mathbb{P}(X\geq k)-\mathbb{P}(Y\geq k)|\leq\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}}),$
@@
 $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\;\geq\;\sup_{k\in\mathbb{Z}}g(k).$ (11)
-Finally, $J=\sum_{k\in\mathbb{Z}}(k-m_{\textup{{\bf p}}})^{2}g(k)$ is defined as in Lemma 7. The technical core of the argument hinges on upper bounding $J$:
-Lemma 8.
-Let $X,Y,\sigma_{\textup{{\bf p}}}^{2}=\operatorname{Var}(X),g,m_{\textup{{\bf p}}},\Delta,J$ be defined as above. Then
+Finally, $J=\sum_{k\in\mathbb{Z}}(k-\mu_{\textup{{\bf p}}})^{2}g(k)$ is defined as in Lemma 2.2. The technical core of the argument hinges on upper bounding $J$:
+Lemma 2.3.
+Let $X,Y,\sigma_{\textup{{\bf p}}}^{2}=\operatorname{Var}(X),g,\mu_{\textup{{\bf p}}},\Delta,J$ be defined as above. Then
 $J\;\leq\;2\,\Delta\,\bigl(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}\bigr).$
@@
 $g(k)=\mathbb{P}(X\geq k)-\mathbb{P}(Y\geq k)=\int_{0}^{1}\sum_{i=1}^{n}\Delta_{i}\,\mathbb{P}\bigl(T_{i}(t)=k-1\bigr)\,\mathrm{d}t,\qquad k\in\mathbb{Z}.$
-In this formulation, the functional $J=\sum_{k}(k-m_{\textup{{\bf p}}})^{2}g(k)$ becomes
-$J$ $=\int_{0}^{1}\sum_{i=1}^{n}\Delta_{i}\sum_{k\in\mathbb{Z}}(k-m_{\textup{{\bf p}}})^{2}\mathbb{P}\bigl(T_{i}(t)=k-1\bigr)\,\mathrm{d}t.$
+In this formulation, the functional $J=\sum_{k}(k-\mu_{\textup{{\bf p}}})^{2}g(k)$ becomes
+$J$ $=\int_{0}^{1}\sum_{i=1}^{n}\Delta_{i}\sum_{k\in\mathbb{Z}}(k-\mu_{\textup{{\bf p}}})^{2}\mathbb{P}\bigl(T_{i}(t)=k-1\bigr)\,\mathrm{d}t.$
 Changing variable j=k1j=k-1,
-$\sum_{k\in\mathbb{Z}}(k-m_{\textup{{\bf p}}})^{2}\mathbb{P}\bigl(T_{i}(t)=k-1\bigr)=\mathbb{E}_{t}\bigl[(T_{i}(t)+1-m_{\textup{{\bf p}}})^{2}\bigr],$
+$\sum_{k\in\mathbb{Z}}(k-\mu_{\textup{{\bf p}}})^{2}\mathbb{P}\bigl(T_{i}(t)=k-1\bigr)=\mathbb{E}_{t}\bigl[(T_{i}(t)+1-\mu_{\textup{{\bf p}}})^{2}\bigr],$
 where $\mathbb{E}_{t}$ denotes expectation under the product law with parameters $\{r_{j}(t)\}_{j}$. Hence
-$J=\int_{0}^{1}\sum_{i=1}^{n}\Delta_{i}\,\mathbb{E}_{t}\bigl[(T_{i}(t)+1-m_{\textup{{\bf p}}})^{2}\bigr]\,\mathrm{d}t.$ (12)
-We bound $\mathbb{E}_{t}[(T_{i}(t)+1-m_{\textup{{\bf p}}})^{2}]$ uniformly in $t,i$ via the decomposition $\mathbb{E}[(Z-\alpha)]^{2}=\operatorname{Var}(Z)+(\mathbb{E}Z-\alpha)^{2}$, where $Z=T_{i}(t)+1$ and $\alpha=m_{\textup{{\bf p}}}$.
+$J=\int_{0}^{1}\sum_{i=1}^{n}\Delta_{i}\,\mathbb{E}_{t}\bigl[(T_{i}(t)+1-\mu_{\textup{{\bf p}}})^{2}\bigr]\,\mathrm{d}t.$ (12)
+We bound $\mathbb{E}_{t}[(T_{i}(t)+1-\mu_{\textup{{\bf p}}})^{2}]$ uniformly in $t,i$ via the decomposition $\mathbb{E}[(Z-\alpha)]^{2}=\operatorname{Var}(Z)+(\mathbb{E}Z-\alpha)^{2}$, where $Z=T_{i}(t)+1$ and $\alpha=\mu_{\textup{{\bf p}}}$.
 To bound the variance term, note that f(x)=x(1x)f(x)=x(1-x) is $1$-Lipschitz on $[0,1]$, and so
@@
 To bound the bias term, we have
-$\mathbb{E}_{t}[S(t)]=m_{\textup{{\bf q}}}+t\Delta,$
-and $m_{\textup{{\bf p}}}=m_{\textup{{\bf q}}}+\Delta$, so $\mathbb{E}_{t}[S(t)]-m_{\textup{{\bf p}}}=-(1-t)\Delta$. Also $r_{i}(t)=q_{i}+t\Delta_{i}$, so
-$\mathbb{E}_{t}[T_{i}(t)]=m_{\textup{{\bf q}}}+t\Delta-(q_{i}+t\Delta_{i}),$
+$\mathbb{E}_{t}[S(t)]=\mu_{\textup{{\bf q}}}+t\Delta,$
+and $\mu_{\textup{{\bf p}}}=\mu_{\textup{{\bf q}}}+\Delta$, so $\mathbb{E}_{t}[S(t)]-\mu_{\textup{{\bf p}}}=-(1-t)\Delta$. Also $r_{i}(t)=q_{i}+t\Delta_{i}$, so
+$\mathbb{E}_{t}[T_{i}(t)]=\mu_{\textup{{\bf q}}}+t\Delta-(q_{i}+t\Delta_{i}),$
 and hence
-$M_{i}(t):=\mathbb{E}_{t}[T_{i}(t)+1-m_{\textup{{\bf p}}}]=1-q_{i}-\Delta+t(\Delta-\Delta_{i}).$
+$M_{i}(t):=\mathbb{E}_{t}[T_{i}(t)+1-\mu_{\textup{{\bf p}}}]=1-q_{i}-\Delta+t(\Delta-\Delta_{i}).$
 It is straightforward to see that $|M_{i}(t)|\leq\max(1,\Delta)$ and thus,
-$\mathbb{E}_{t}\bigl[(T_{i}(t)+1-m_{\textup{{\bf p}}})^{2}\bigr]$ $=$ $\operatorname{Var}_{t}(T_{i}(t))+M_{i}(t)^{2}$
+$\mathbb{E}_{t}\bigl[(T_{i}(t)+1-\mu_{\textup{{\bf p}}})^{2}\bigr]$ $=$ $\operatorname{Var}_{t}(T_{i}(t))+M_{i}(t)^{2}$
 $\leq$ $(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2})+(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2})=2(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}).$
@@
 as claimed. ∎
-Corollary 9 (Proof of Theorem 2).
+Corollary 2.4 (Proof of Theorem 1.2).
 Let $(\textup{{\bf p}},\textup{{\bf q}})$ be a dominating pair with laws $S_{\textup{{\bf p}}},S_{\textup{{\bf q}}}$, sums $S_{\textup{{\bf p}}},S_{\textup{{\bf q}}}$, $\sigma_{\textup{{\bf p}}}^{2}:=\operatorname{Var}(S_{\textup{{\bf p}}})$, and $\Delta:=\mathbb{E}S_{\textup{{\bf p}}}-\mathbb{E}S_{\textup{{\bf q}}}$.
@@
 $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\;\geq\;c\,\min\Bigl(1,\frac{\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}}\Bigr),$
-for some universal constant $c\geq\frac{1}{12}$.
+for some universal constant $c\geq\frac{1}{9}$.
 Proof.
 Recall $g(k):=\mathbb{P}(S_{\textup{{\bf p}}}\geq k)-\mathbb{P}(S_{\textup{{\bf q}}}\geq k)\geq 0$ and
-$\Delta=\sum_{k}g(k),\qquad J=\sum_{k}(k-m_{\textup{{\bf p}}})^{2}g(k),\qquad m_{\textup{{\bf p}}}=\mathbb{E}S_{\textup{{\bf p}}}.$
-By (11), $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\geq\sup_{k\in\mathbb{Z}}g(k)$. By Lemma 8,
+$\Delta=\sum_{k}g(k),\qquad J=\sum_{k}(k-\mu_{\textup{{\bf p}}})^{2}g(k),\qquad\mu_{\textup{{\bf p}}}=\mathbb{E}S_{\textup{{\bf p}}}.$
+By (11), $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\geq\sup_{k\in\mathbb{Z}}g(k)$. By Lemma 2.3,
 $J\;\leq\;2\,\Delta\,(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}).$
-Apply Lemma 7 with $G=\Delta$ to obtain
-$\sup_{k}g(k)\;\geq\;\frac{3\Delta}{16\sqrt{2(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2})}+4}.$
+Apply Lemma 2.2 with $G=\Delta$ to obtain
+$\sup_{k}g(k)\;\geq\;\frac{\Delta}{4\sqrt{2(\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2})}+1}.$
 We compare $\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}}$ with $\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}$ and consider the two cases.
-If $\Delta\leq\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}$, then $\Delta^{2}\leq\sigma_{\textup{{\bf p}}}^{2}+1$, so
-$\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}\leq 2(\sigma_{\textup{{\bf p}}}^{2}+1),$
-and
-$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\geq\frac{3\Delta}{16\sqrt{2}\sqrt{2(\sigma_{\textup{{\bf p}}}^{2}+1)}+4}=\frac{3\Delta}{32\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+4}$
-$\geq\frac{3\Delta}{36\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}}=\frac{\Delta}{12\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}}.$
-If $\Delta\geq\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}$, then $\Delta^{2}\geq\sigma_{\textup{{\bf p}}}^{2}+1$, so
-$\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}\leq 2\Delta^{2},$
-and
-$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\geq\frac{3\Delta}{16\sqrt{2}\sqrt{2\Delta^{2}}+4}=\frac{3\Delta}{32\Delta+4}\geq\frac{3\Delta}{36\Delta}=\frac{1}{12}.$
+If $\Delta\leq\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}$, then $\Delta^{2}\leq\sigma_{\textup{{\bf p}}}^{2}+1$, so $\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}\leq 2(\sigma_{\textup{{\bf p}}}^{2}+1),$ and
+$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\geq\frac{\Delta}{4\sqrt{2}\sqrt{2(\sigma_{\textup{{\bf p}}}^{2}+1)}+1}=\frac{\Delta}{8\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+1}\geq\frac{\Delta}{9\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}}.$
+If $\Delta\geq\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}$, then $\Delta^{2}\geq\sigma_{\textup{{\bf p}}}^{2}+1$, so $\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}\leq 2\Delta^{2},$ and
+$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\geq\frac{\Delta}{4\sqrt{2}\sqrt{2\Delta^{2}}+1}=\frac{\Delta}{8\Delta+1}\geq\frac{1}{9}.$
 Combining the two cases proves the claim. ∎
-2.3 Proof of Theorem 3
+2.3 Proof of Theorem 1.3
 Both inequalities invoke the data processing inequality; the first in the form
@@
 and the second in the form $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$.
-Given these, the claim is an immediate consequence of Theorem 2.
-2.4 Proof of Lemma 4
+Given these, the claim is an immediate consequence of Theorem 1.2.
+2.4 Proof of Lemma 1.4
 Proof.
@@
 Define
-$\Delta_{I}(m):=\operatorname{TV}(\operatorname{Bin}(m,\bar{p}_{I}),\operatorname{Bin}(m,\bar{q}_{I})),\qquad\Delta_{J}(k):=\operatorname{TV}(\operatorname{Bin}(k,\bar{p}_{J}),\operatorname{Bin}(k,\bar{q}_{J})).$
+$\tau_{I}(m):=\operatorname{TV}(\operatorname{Bin}(m,\bar{p}_{I}),\operatorname{Bin}(m,\bar{q}_{I})),\qquad\tau_{J}(k):=\operatorname{TV}(\operatorname{Bin}(k,\bar{p}_{J}),\operatorname{Bin}(k,\bar{q}_{J})).$
 Then
-$\delta_{N}\leq\mathbb{E}[\Delta_{I}(M)]+\mathbb{E}[\Delta_{J}(n-M)].$
+$\delta_{N}\leq\mathbb{E}[\tau_{I}(M)]+\mathbb{E}[\tau_{J}(n-M)].$
 It remains to estimate the two expectations. Fix $\theta,\theta^{\prime}\in[0,1]$ and define $f:\mathbb{N}\to\mathbb{R}$ by
@@
 $\mathbb{E}[f(M)]\leq f(m_{0})\,\mathbb{E}\!\left[\left\lceil\frac{M}{m_{0}}\right\rceil\right]\leq f(m_{0})\,\mathbb{E}\!\left[\frac{M}{m_{0}}+1\right]=2f(m_{0}).$
-Apply this with $f=\Delta_{I}$ and $m_{0}=|I|$ (note $\mathbb{E}M=nw=|I|$) to get
-$\mathbb{E}[\Delta_{I}(M)]\leq 2\Delta_{I}(|I|)=2\delta_{I},$
-and with $f=\Delta_{J}$ and $m_{0}=|J|$ (since $\mathbb{E}(n-M)=|J|$) to get
-$\mathbb{E}[\Delta_{J}(n-M)]\leq 2\Delta_{J}(|J|)=2\delta_{J}.$
+Apply this with $f=\tau_{I}$ and $m_{0}=|I|$ (note $\mathbb{E}M=nw=|I|$) to get
+$\mathbb{E}[\tau_{I}(M)]\leq 2\tau_{I}(|I|)=2\delta_{I},$
+and with $f=\tau_{J}$ and $m_{0}=|J|$ (since $\mathbb{E}(n-M)=|J|$) to get
+$\mathbb{E}[\tau_{J}(n-M)]\leq 2\tau_{J}(|J|)=2\delta_{J}.$
 This completes the proof. ∎
-Acknowledgments.
-Some of the proofs were assisted by ChatGPT Pro; we thank Sébastien Bubeck for making this version available to us.
-We also thank Mark Sellke and Nikita Zhivotovskiy for pointing out errors in earlier versions.
+Acknowledgements
+Some of the proofs were assisted by ChatGPT Pro.
+We thank Sébastien Bubeck, Mark Sellke and Nikita Zhivotovskiy for pointing out errors in earlier versions.
+We are grateful to Lutz Mattner and Bero Roos for detailed and insightful comments on the manuscript.
+Appendix A Homogenization is a Markov Kernel in the symmetric case
+We show that for symmetric $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ — i.e., those satisfying $\textup{{\bf p}}=\textup{{\bf 1}}-\textup{{\bf q}}$, homogenization can be realized by a Markov kernel.
+We first consider the case $n=2$:
+Lemma A.1 (averaging centered Bernoullis).
+Let $\alpha,\beta\in(0,1)$ and put $\gamma=(\alpha+\beta)/2$. Then there exists a randomized function $F_{\alpha,\beta}:\left\{0,1\right\}^{2}\to\left\{0,1\right\}^{2}$ with the following property:
+$X\sim\operatorname{Ber}\left(\frac{1}{2}+\frac{\alpha}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}+\frac{\beta}{2}\right)$ $\implies$ $F_{\alpha,\beta}(X)\sim\operatorname{Ber}\left(\frac{1}{2}+\frac{\gamma}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}+\frac{\gamma}{2}\right),$
+$Y\sim\operatorname{Ber}\left(\frac{1}{2}-\frac{\alpha}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}-\frac{\beta}{2}\right)$ $\implies$ $F_{\alpha,\beta}(Y)\sim\operatorname{Ber}\left(\frac{1}{2}-\frac{\gamma}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}-\frac{\gamma}{2}\right).$
+Proof.
+As in the proof of [9, Lemma 2.3], it suffices to write down the stochastic vectors representing $X,Y$ and $\hat{X}\sim\operatorname{Ber}\left(\frac{1}{2}+\frac{\gamma}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}+\frac{\gamma}{2}\right),\hat{Y}\sim\operatorname{Ber}\left(\frac{1}{2}-\frac{\gamma}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}-\frac{\gamma}{2}\right)$, the stochastic matrix $M$ representing $F_{\alpha,\beta}$, and to verify that $\hat{X}=XM$ and $\hat{Y}=YM$.
+For readability, we mildly abuse notation and blur the distinction between the random variable $Z\sim\operatorname{Ber}(p)\otimes\operatorname{Ber}(q)$ and its stochastic vector:
+$Z=\left(\mathbb{P}(Z=00),\mathbb{P}(Z=01),\mathbb{P}(Z=10),\mathbb{P}(Z=11)\right)$ $=$ $\left((1-p)(1-q),(1-p)q,p(1-q),pq\right).$
+With this convention, the four stochastic vectors are:
+$X$ $=$ $\left(\left(\frac{1}{2}-\frac{\alpha}{2}\right)\left(\frac{1}{2}-\frac{\beta}{2}\right),\left(\frac{1}{2}-\frac{\alpha}{2}\right)\left(\frac{1}{2}+\frac{\beta}{2}\right),\left(\frac{1}{2}+\frac{\alpha}{2}\right)\left(\frac{1}{2}-\frac{\beta}{2}\right),\left(\frac{1}{2}+\frac{\alpha}{2}\right)\left(\frac{1}{2}+\frac{\beta}{2}\right)\right),$
+$Y$ $=$ $\left(\left(\frac{1}{2}+\frac{\alpha}{2}\right)\left(\frac{1}{2}+\frac{\beta}{2}\right),\left(\frac{1}{2}+\frac{\alpha}{2}\right)\left(\frac{1}{2}-\frac{\beta}{2}\right),\left(\frac{1}{2}-\frac{\alpha}{2}\right)\left(\frac{1}{2}+\frac{\beta}{2}\right),\left(\frac{1}{2}-\frac{\alpha}{2}\right)\left(\frac{1}{2}-\frac{\beta}{2}\right)\right),$
+$\hat{X}$ $=$ $\left(\left(\frac{1}{2}-\frac{\gamma}{2}\right)\left(\frac{1}{2}-\frac{\gamma}{2}\right),\left(\frac{1}{2}-\frac{\gamma}{2}\right)\left(\frac{1}{2}+\frac{\gamma}{2}\right),\left(\frac{1}{2}+\frac{\gamma}{2}\right)\left(\frac{1}{2}-\frac{\gamma}{2}\right),\left(\frac{1}{2}+\frac{\gamma}{2}\right)\left(\frac{1}{2}+\frac{\gamma}{2}\right)\right),$
+$\hat{Y}$ $=$ $\left(\left(\frac{1}{2}+\frac{\gamma}{2}\right)\left(\frac{1}{2}+\frac{\gamma}{2}\right),\left(\frac{1}{2}+\frac{\gamma}{2}\right)\left(\frac{1}{2}-\frac{\gamma}{2}\right),\left(\frac{1}{2}-\frac{\gamma}{2}\right)\left(\frac{1}{2}+\frac{\gamma}{2}\right),\left(\frac{1}{2}-\frac{\gamma}{2}\right)\left(\frac{1}{2}-\frac{\gamma}{2}\right)\right).$
+With $\eta:=\frac{(\alpha-\beta)^{2}}{8(1-\alpha\beta)}$, we define the $4\times 4$ row-stochastic matrix $M$:
+$M=\left(\begin{array}[]{cccc}1&0&0&0\\ \eta&\frac{1}{2}-\eta&\frac{1}{2}-\eta&\eta\\ \eta&\frac{1}{2}-\eta&\frac{1}{2}-\eta&\eta\\ 0&0&0&1\end{array}\right).$ (19)
+Verifying the identities $\hat{X}=XM$ and $\hat{Y}=YM$ is tedious but quite straightforward.
+Obviously, the rows of $M$ sum to $1$.
+It only remains to show that the entries of $M$ are nonnegative, or equivalently, that $\eta=\frac{(\alpha-\beta)^{2}}{8(1-\alpha\beta)}\leq 1/2$.
+The latter is equivalent to ${\alpha^{2}-2\alpha\beta+\beta^{2}}\leq{4(1-\alpha\beta)}$.
+Expanding and collecting like terms yields the equivalent claim $4-(\alpha+\beta)^{2}\geq 0$, which obviously holds for $\alpha,\beta\in(0,1)$.
+∎
+For general $n$, the pairwise averaging kernel $M=M(\alpha_{i},\beta_{i})$ defined in (19) can be applied repeatedly to the pairs $(i,j)$.
+Although the kernel itself changes with each iteration, a standard compactness argument shows that for any symmetric pair $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ there exists a Markov kernel $M_{\textup{{\bf p}},\textup{{\bf q}}}$ simultaneously achieving $M_{\textup{{\bf p}},\textup{{\bf q}}}:\operatorname{Ber}(\textup{{\bf p}})\mapsto\operatorname{Ber}(\bar{p})^{\otimes n}$ and $M_{\textup{{\bf p}},\textup{{\bf q}}}:\operatorname{Ber}(\textup{{\bf q}})\mapsto\operatorname{Ber}(\bar{q})^{\otimes n}$.

### v3

@@
 TV homogenization inequalities
-Abstract
+Abstract.
 We study the total variation distance under two information-erasing maps on inhomogeneous Bernoulli product measures: summation and homogenization.
@@
 The argument is based on an explicit two-sided control of the TV distance between Poisson binomials, obtained via a parameter interpolation and a second-moment extraction.
-1 Introduction
+Key words and phrases:
+total variation; homogenization; Poisson binomial; Bernoulli product measures
+2020 Mathematics Subject Classification:
+60E15, 60F05
+1. Introduction
 For $\textup{{\bf p}}=(p_{1},\dots,p_{n})\in[0,1]^{n}$, the Poisson–binomial law, introduced by [16] [^1] See [17] for a modern English translation.
@@
 Equivalently, if $X_{i}\sim\operatorname{Ber}(p_{i})$ are independent, then $S_{\textup{{\bf p}}}:=\sum_{i=1}^{n}X_{i}$ is distributed according to this law.
-The sum map is easily seen to be a Markov kernel, and hence the Data Processing Inequality [16, Theorem 7.4] applies: $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$, where $\operatorname{TV}$ is the total variation distance.
+The sum map is easily seen to be a Markov kernel, and hence the Data Processing Inequality [18, Theorem 7.4] applies: $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$, where $\operatorname{TV}$ is the total variation distance.
 Intuitively, since information is lost under the sum map, the distributions can only become closer to each other.
@@
 However, $\mathbb{P}(Z=2)=\binom{n}{2}\left(\frac{t}{n}\right)^{2}\left(1-\frac{t}{n}\right)^{n-2}$, which is not affine in tt.
-[^2] For the special case of $\textup{{\bf p}},\textup{{\bf q}}$ that are symmetric about $\frac{1}{2}$, meaning that $\textup{{\bf q}}=\textup{{\bf 1}}-\textup{{\bf p}}$, it turns out that the homogenization operator can be realized as a Markov kernel; see Appendix.
+[^2] For the special case of complementary $\textup{{\bf p}},\textup{{\bf q}}$, meaning that $\textup{{\bf q}}=\textup{{\bf 1}}-\textup{{\bf p}}$, it turns out that the homogenization operator can be realized as a Markov kernel; see Appendix.
 This indicates that the standard information-theoretic argument for proving data processing inequalities does not apply, necessitating novel tools.
@@
 For $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, define the quantities
-$\Delta:=\sum_{i=1}^{n}|p_{i}-q_{i}|,\quad\sigma_{\textup{{\bf p}}}^{2}:=\operatorname{Var}(S_{\textup{{\bf p}}})=\sum_{i=1}^{n}p_{i}(1-p_{i}),\quad\Phi(\textup{{\bf p}},\textup{{\bf q}}):=\min\!\left(1,\frac{\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}}\right).$ (1)
+(1) $\Delta:=\sum_{i=1}^{n}|p_{i}-q_{i}|,\quad\sigma_{\textup{{\bf p}}}^{2}:=\operatorname{Var}(S_{\textup{{\bf p}}})=\sum_{i=1}^{n}p_{i}(1-p_{i}),\quad\Phi(\textup{{\bf p}},\textup{{\bf q}}):=\min\!\left(1,\frac{\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}}\right).$
 We discover a simple analytic approximation of $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ for dominating pairs $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, $p_{i}\geq q_{i}$, up to universal multiplicative constants:
-$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\asymp\Phi(\textup{{\bf p}},\textup{{\bf q}}),\qquad\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n},\textup{{\bf p}}\geq\textup{{\bf q}}.$ (2)
+(2) $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\asymp\Phi(\textup{{\bf p}},\textup{{\bf q}}),\qquad\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n},\textup{{\bf p}}\geq\textup{{\bf q}}.$
 For general $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, only the upper bound $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\lesssim\Phi(\textup{{\bf p}},\textup{{\bf q}})$ holds, which can be sharpened to $\min(\Phi(\textup{{\bf p}},\textup{{\bf q}}),\Phi(\textup{{\bf q}},\textup{{\bf p}}))$ by symmetry.
-Unlike $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$, whose exact computation is infeasible (see below), $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ is computable exactly in time $O(n^{2})$ via a simple recurrence [20, Section 5].
+Unlike $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$, whose exact computation is infeasible (see below), $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ is computable exactly in time $O(n^{2})$ via a simple recurrence [22, Section 5].
 Still, one should not expect any analytically tractable analogue of (2) for general $\textup{{\bf p}},\textup{{\bf q}}$ due to cancellations and convolutional smoothing.
 As an application of these results, we prove the TV homogenization inequality for all $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$:
-$\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\gtrsim$ $\operatorname{TV}(\operatorname{Ber}(\bar{p})^{\otimes n},\operatorname{Ber}(\bar{q})^{\otimes n})=\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q})).$ (3)
-Notation.
-We write $n]:=\left\{1,\ldots,n\right\}$. For two distributions $P,Q$ on a finite set $\Omega$, their total variation distance is defined by
-$\operatorname{TV}(P,Q)$ $=$ $\frac{1}{2}\sum_{\omega\in\Omega}|P(\omega)-Q(\omega)|.$ (4)
+(3) $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\gtrsim$ $\operatorname{TV}(\operatorname{Ber}(\bar{p})^{\otimes n},\operatorname{Ber}(\bar{q})^{\otimes n})=\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q})).$
+Notation. We write $n]:=\left\{1,\ldots,n\right\}$. For two distributions $P,Q$ on a finite set $\Omega$, their total variation distance is defined by
+(4) $\operatorname{TV}(P,Q)$ $=$ $\frac{1}{2}\sum_{\omega\in\Omega}|P(\omega)-Q(\omega)|.$
 If $X,Y$ are random variables with laws $P,Q$, respectively, then we write $\operatorname{TV}(X,Y):=\operatorname{TV}(P,Q)$.
@@
 The notation in (1) will persist throughout the paper.
-Main results.
-Our first result implies the upper bound in (2). Let $\eta_{\mathrm{BCV}}$ denote the sharp universal constant from [1, Theorem 1], and set
-$C_{\mathrm{BCV}}:=\sqrt{\frac{5}{4}+\eta_{\mathrm{BCV}}^{2}}.$
-Numerically (see [1, Remark 1]), $\eta_{\mathrm{BCV}}\approx 0.4688223555$ and hence $C_{\mathrm{BCV}}\approx 1.2123507747$.
+Main results. Our first result implies the upper bound in (2). Let $\eta_{0}$ denote the sharp universal constant from [2, Theorem 1], and set
+$C_{0}:=\sqrt{\frac{5}{4}+\eta_{0}^{2}}.$
+Numerically (see [2, Remark 1]), $\eta_{0}\approx 0.4688223555$ and hence $C_{0}\approx 1.2123507747$.
 Theorem 1.1.
 For $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ and $\Phi$ as in (1),
-$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\leq$ $2C_{\mathrm{BCV}}\,\min(\Phi(\textup{{\bf p}},\textup{{\bf q}}),\Phi(\textup{{\bf q}},\textup{{\bf p}})).$
+$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\leq$ $2C_{0}\,\min(\Phi(\textup{{\bf p}},\textup{{\bf q}}),\Phi(\textup{{\bf q}},\textup{{\bf p}})).$
+L. Mattner [14] has brought to our attention the result of [23], which upper bounds $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ by
+$\frac{5|\mathbb{E}S_{\textup{{\bf p}}}-\mathbb{E}S_{\textup{{\bf q}}}|}{\sqrt{\sigma^{2}_{\textup{{\bf p}}}+\sigma^{2}_{\textup{{\bf q}}}}}+\frac{12|\mathbb{E}S_{\textup{{\bf p}}}-\mathbb{E}S_{\textup{{\bf q}}}|}{{1+\sigma^{2}_{\textup{{\bf p}}}+\sigma^{2}_{\textup{{\bf q}}}}}.$
+This estimate has the attractive property of being topologically equivalent to $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ in the sense that the two converge to 0 on precisely the same set of sequence pairs.
+It is currently an open question whether a version of our $\Phi$ functional could be made topologically equivalent to $\operatorname{TV}$ as well.
 The technical core of the paper is a matching lower bound for dominating pairs:
@@
 $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\geq$ $\frac{1}{9}\Phi(\textup{{\bf p}},\textup{{\bf q}}).$
-The assumption $\textup{{\bf p}}\geq\textup{{\bf q}}$ is essential for the lower bound, as illustrated by the example $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, for even $n$, given by
-$\textup{{\bf p}}=\Bigl(\tfrac{1}{2},\tfrac{1}{2},\dots,\tfrac{1}{2}\Bigr)$, $\textup{{\bf q}}=\Bigl(\tfrac{1}{2}+\delta,\dots,\tfrac{1}{2}+\delta,\ \tfrac{1}{2}-\delta,\dots,\tfrac{1}{2}-\delta\Bigr),$
-with $n/2$ copies of $\tfrac{1}{2}+\delta$ followed by $n/2$ copies of $\tfrac{1}{2}-\delta$.
-The choice
-$\delta=1/\sqrt{n}$ yields $\Phi(\textup{{\bf p}},\textup{{\bf q}})=1$, but $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\asymp 1/n$.
+The assumption $\textup{{\bf p}}\geq\textup{{\bf q}}$ is essential for the lower bound, as illustrated by the example $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, for even $n$, given by $\textup{{\bf p}}=\Bigl(\tfrac{1}{2},\tfrac{1}{2},\dots,\tfrac{1}{2}\Bigr)$, $\textup{{\bf q}}=\Bigl(\tfrac{1}{2}+\delta,\dots,\tfrac{1}{2}+\delta,\ \tfrac{1}{2}-\delta,\dots,\tfrac{1}{2}-\delta\Bigr),$ with $n/2$ copies of $\tfrac{1}{2}+\delta$ followed by $n/2$ copies of $\tfrac{1}{2}-\delta$.
+The choice $\delta=1/\sqrt{n}$ yields $\Phi(\textup{{\bf p}},\textup{{\bf q}})=1$, but $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\asymp 1/n$.
 In general, for any $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ and any partition $I,J$ of $[n]$, the trivial estimate via the partitioned pairs $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\leq\operatorname{TV}(S_{\textup{{\bf p}}_{I}},S_{\textup{{\bf q}}_{I}})+\operatorname{TV}(S_{\textup{{\bf p}}_{J}},S_{\textup{{\bf q}}_{J}})$ holds.
@@
 $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\geq$ $c\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q})),$
-where $c\geq\frac{1}{72C_{\mathrm{BCV}}}\approx 0.0115$ is a universal constant.
+where $c\geq\frac{1}{72C_{0}}\approx 0.0115$ is a universal constant.
 Remark.
-As the remark preceding Theorem 2.1 indicates, a sharpened constant $c\geq\frac{1}{36C_{\mathrm{BCV}}}$ is straightforward to obtain via analogous arguments.
+As the remark preceding Theorem 2.1 indicates, a sharpened constant $c\geq\frac{1}{36C_{0}}$ is straightforward to obtain via analogous arguments.
 The choice $\textup{{\bf p}}=(1-2\varepsilon,\tfrac{1}{2})$, $\textup{{\bf q}}=(1,\tfrac{1}{2}+\varepsilon)$, $\varepsilon\to 0$ demonstrates that it is possible for homogenization to (slightly) increase the TV distance, and also that $8/9$ is an upper bound on $c$.
@@
 $\Phi(\textup{{\bf p}}_{I},\textup{{\bf q}}_{I})+\Phi(\textup{{\bf q}}_{J},\textup{{\bf p}}_{J})$ $\geq$ $\Phi(\bar{p}_{I}\boldsymbol{1},\bar{q}_{I}\boldsymbol{1})+\Phi(\bar{q}_{J}\boldsymbol{1},\bar{p}_{J}\boldsymbol{1})$
-$\geq$ $\frac{1}{2C_{\mathrm{BCV}}}\left(\operatorname{TV}(\operatorname{Bin}(|I|,\bar{p}_{I}),\operatorname{Bin}(|I|,\bar{q}_{I}))+\operatorname{TV}(\operatorname{Bin}(|J|,\bar{p}_{J}),\operatorname{Bin}(|J|,\bar{q}_{J}))\right)$
-$\geq$ $\frac{1}{4C_{\mathrm{BCV}}}\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q})),$
+$\geq$ $\frac{1}{2C_{0}}\left(\operatorname{TV}(\operatorname{Bin}(|I|,\bar{p}_{I}),\operatorname{Bin}(|I|,\bar{q}_{I}))+\operatorname{TV}(\operatorname{Bin}(|J|,\bar{p}_{J}),\operatorname{Bin}(|J|,\bar{q}_{J}))\right)$
+$\geq$ $\frac{1}{4C_{0}}\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q})),$
 where the last inequality is by Lemma 1.4.
@@
 Related work.
-This paper follows a program initiated by [9] to provide simple, analytically tractable upper and lower estimates on $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ in terms of the $\textup{{\bf p}},\textup{{\bf q}}$; the latter’s contribution consisted of sharpening of the trivial lower bound $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{\infty}$ to $\gtrsim\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{2}$.
+This paper follows a program initiated by [11] to provide simple, analytically tractable upper and lower estimates on $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ in terms of the $\textup{{\bf p}},\textup{{\bf q}}$; the latter’s contribution sharpened the trivial lower bound $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{\infty}$ to $\gtrsim\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{2}$.
 We note that generally, this result is incomparable with the lower bound $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\gtrsim$ $\Phi(\textup{{\bf p}},\textup{{\bf q}})$ implied by Theorem 1.2 for $\textup{{\bf p}}\geq\textup{{\bf q}}$.
@@
 A delicate structural result due to [6] (also ultimately yielding an additive estimate) shows that every Poisson binomial admits an $\varepsilon$-approximation in TV by a distribution of one of two canonical “compressed” types: either a sparse $S_{\textup{{\bf p}}}$ with only $O(1/\varepsilon^{3})$ nontrivial summands, or a near-binomial form.
-Of particular relevance to this work is [1], whose structural results we build upon in Theorem 1.1.
+Of particular relevance to this work is [2, 3], whose structural results we build upon in Theorem 1.1.
 More generally, $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ admits classic approximations in terms of more analytically tractable proxies such as KL-divergence and the Hellinger distance; their properties and limitations are discussed in [11].
@@
 An efficient randomized algorithm guaranteeing a $1\pm\varepsilon$ multiplicative approximation with confidence $1-\delta$, in time $O(\frac{n^{2}}{\varepsilon^{2}}\log\frac{1}{\delta})$ was discovered by [8], and later derandomized by [9].
-2 Proofs
-2.1 Proof of Theorem 1.1
+2. Proofs
+2.1. Proof of Theorem 1.1
 We will prove a more general, symmetric bound, which immediately implies the one in Theorem 1.1.
-The constant $2C_{\mathrm{BCV}}$ in the latter’s bound could be sharpened by a factor of $2$ via a direct (asymmetric) proof.
-L.
-Mattner [12] has brought to our attention the result of [21], which upper bounds $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ by
-$\frac{5|\mathbb{E}S_{\textup{{\bf p}}}-\mathbb{E}S_{\textup{{\bf q}}}|}{\sqrt{\sigma^{2}_{\textup{{\bf p}}}+\sigma^{2}_{\textup{{\bf q}}}}}+\frac{12|\mathbb{E}S_{\textup{{\bf p}}}-\mathbb{E}S_{\textup{{\bf q}}}|}{{1+\sigma^{2}_{\textup{{\bf p}}}+\sigma^{2}_{\textup{{\bf q}}}}}.$
-This estimate has the attractive property of being topologically equivalent to $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ in the sense that the two converge to 0 on precisely the same set of sequence pairs.
-It is currently an open question whether a version of our $\Phi$ functional could be made topologically equivalent to $\operatorname{TV}$ as well.
+The constant $2C_{0}$ in the latter’s bound could be sharpened by a factor of $2$ via a direct (asymmetric) proof.
 Theorem 2.1.
 For $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ and $S_{\textup{{\bf p}}},S_{\textup{{\bf q}}}$ the corresponding Poisson binomials with variances $\sigma_{\textup{{\bf p}}}^{2}$, $\sigma_{\textup{{\bf q}}}^{2}$, we have
-$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\leq$ $\frac{2C_{\mathrm{BCV}}\,\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+\sqrt{\sigma_{\textup{{\bf q}}}^{2}+1}}.$
+$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\leq$ $\frac{2C_{0}\,\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+\sqrt{\sigma_{\textup{{\bf q}}}^{2}+1}}.$
 Proof.
@@
 Shift-TV for unimodal pmfs. If $Z$ is integer-valued with unimodal pmf $h(k)=\mathbb{P}(Z=k)$, then
-$\operatorname{TV}(Z,Z+1)=\max_{k}h(k),$ (5)
+(5) $\operatorname{TV}(Z,Z+1)=\max_{k}h(k),$
 by a simple telescoping argument.
-[^3] M.
-Lutz [12] informs us that this result is contained in [2] with a slightly worse constant.
-Anti-concentration for Poisson binomials (Baillon–Cominetti–Vaisman).
-If $Z=\sum_{j=1}^{m}\operatorname{Ber}(r_{j})$ has variance $v=\operatorname{Var}(Z)$, then [1, Theorem 1] proves the sharp bound
-$\max_{k}\mathbb{P}(Z=k)\leq\frac{\eta_{\mathrm{BCV}}}{\sqrt{v}}\qquad(v>0),$ (6)
-where $\eta_{\mathrm{BCV}}$ is a universal constant ($\eta_{\mathrm{BCV}}\approx 0.4688223555$, [1, Remark 1]). In particular, since $\max_{k}\mathbb{P}(Z=k)\leq 1$ always, we have for all $v\geq 0$
-$\max_{k}\mathbb{P}(Z=k)\leq\min\!\left\{1,\ \frac{\eta_{\mathrm{BCV}}}{\sqrt{v}}\right\}.$ (7)
-Unimodality of Poisson binomial pmfs. It is a classic consequence of the Aissen-Edrei-Schoenberg-Whitney theorem [20, Section 4] that the Poisson binomial
-270]
-has a log-concave, and hence unimodal, law. To combine these three ingredients, we begin with the interpolation. For $t\in[0,1]$ define
+Anti-concentration for Poisson binomials (Barbour-Jensen; Baillon–Cominetti–Vaisman).
+If $Z=\sum_{j=1}^{m}\operatorname{Ber}(r_{j})$ has variance $v=\operatorname{Var}(Z)$, then [2, Theorem 1] proves the sharp bound
+(6) $\max_{k}\mathbb{P}(Z=k)\leq\frac{\eta_{0}}{\sqrt{v}}\qquad(v>0),$
+where $\eta_{0}$ is a universal constant ($\eta_{0}\approx 0.4688223555$, [2, Remark 1]). [^3] L. Mattner [14] informs us that this result is contained in [3] with a slightly worse constant.
+In particular, since $\max_{k}\mathbb{P}(Z=k)\leq 1$ always, we have for all $v\geq 0$
+(7) $\max_{k}\mathbb{P}(Z=k)\leq\min\!\left\{1,\ \frac{\eta_{0}}{\sqrt{v}}\right\}.$
+Unimodality of Poisson binomial pmfs.
+It is a classic consequence of the Aissen-Edrei-Schoenberg-Whitney theorem [22, Section 4] that the Poisson binomial has a log-concave, and hence unimodal, law.
+To combine these three ingredients, we begin with the interpolation.
+For $t\in[0,1]$ define
 $r_{i}(t):=(1-t)q_{i}+tp_{i},\qquad S(t):=\sum_{i=1}^{n}\operatorname{Ber}(r_{i}(t)),\qquad t\in[0,1],$
@@
 By (7),
-$\max_{k}\mathbb{P}(T_{i}(t)=k)\leq\min\!\left\{1,\ \frac{\eta_{\mathrm{BCV}}}{\sqrt{\operatorname{Var}(T_{i}(t))}}\right\}.$
+$\max_{k}\mathbb{P}(T_{i}(t)=k)\leq\min\!\left\{1,\ \frac{\eta_{0}}{\sqrt{\operatorname{Var}(T_{i}(t))}}\right\}.$
 Since $S(t)=T_{i}(t)+X_{i}(t)$ with independence and $\operatorname{Var}(X_{i}(t))\leq\tfrac{1}{4}$, we have
@@
 Let $x_{+}:=\max\{x,0\}$. Then
-$\max_{k}\mathbb{P}(T_{i}(t)=k)\leq\min\!\left\{1,\ \frac{\eta_{\mathrm{BCV}}}{\sqrt{\bigl(\operatorname{Var}(S(t))-\frac{1}{4}\bigr)_{+}}}\right\}.$
-We now upper bound this “singular” expression by a smooth envelope. Set $C_{\mathrm{BCV}}:=\sqrt{\tfrac{5}{4}+\eta_{\mathrm{BCV}}^{2}}$. Then for every $x\geq 0$ we have
-$\min\!\left\{1,\ \frac{\eta_{\mathrm{BCV}}}{\sqrt{(x-\frac{1}{4})_{+}}}\right\}\leq\frac{C_{\mathrm{BCV}}}{\sqrt{x+1}}.$ (8)
-Indeed, if $x\leq\tfrac{1}{4}+\eta_{\mathrm{BCV}}^{2}$, the left-hand side equals $1$ and the right-hand side is $C_{\mathrm{BCV}}/\sqrt{x+1}\geq C_{\mathrm{BCV}}/\sqrt{\tfrac{5}{4}+\eta_{\mathrm{BCV}}^{2}}=1$.
-If $x\geq\tfrac{1}{4}+\eta_{\mathrm{BCV}}^{2}$, then the left-hand side equals $\eta_{\mathrm{BCV}}/\sqrt{x-\tfrac{1}{4}}$ and squaring shows that (8) is equivalent to $x\geq\tfrac{1}{4}+\eta_{\mathrm{BCV}}^{2}$.
+$\max_{k}\mathbb{P}(T_{i}(t)=k)\leq\min\!\left\{1,\ \frac{\eta_{0}}{\sqrt{\bigl(\operatorname{Var}(S(t))-\frac{1}{4}\bigr)_{+}}}\right\}.$
+We now upper bound this “singular” expression by a smooth envelope. Set $C_{0}:=\sqrt{\tfrac{5}{4}+\eta_{0}^{2}}$. Then for every $x\geq 0$ we have
+(8) $\min\!\left\{1,\ \frac{\eta_{0}}{\sqrt{(x-\frac{1}{4})_{+}}}\right\}\leq\frac{C_{0}}{\sqrt{x+1}}.$
+Indeed, if $x\leq\tfrac{1}{4}+\eta_{0}^{2}$, the left-hand side equals $1$ and the right-hand side is $C_{0}/\sqrt{x+1}\geq C_{0}/\sqrt{\tfrac{5}{4}+\eta_{0}^{2}}=1$.
+If $x\geq\tfrac{1}{4}+\eta_{0}^{2}$, then the left-hand side equals $\eta_{0}/\sqrt{x-\tfrac{1}{4}}$ and squaring shows that (8) is equivalent to $x\geq\tfrac{1}{4}+\eta_{0}^{2}$.
 Applying (8) with $x=\operatorname{Var}(S(t))$ gives
-$\operatorname{TV}\!\bigl(T_{i}(t)+1,T_{i}(t)\bigr)\leq\frac{C_{\mathrm{BCV}}}{\sqrt{\operatorname{Var}(S(t))+1}},$
+$\operatorname{TV}\!\bigl(T_{i}(t)+1,T_{i}(t)\bigr)\leq\frac{C_{0}}{\sqrt{\operatorname{Var}(S(t))+1}},$
 and substituting back,
-$|f_{A}^{\prime}(t)|\leq\frac{C_{\mathrm{BCV}}\,\Delta}{\sqrt{\operatorname{Var}(S(t))+1}}.$ (9)
+(9) $|f_{A}^{\prime}(t)|\leq\frac{C_{0}\,\Delta}{\sqrt{\operatorname{Var}(S(t))+1}}.$
 Since $u\mapsto u(1-u)$ is concave,
@@
 Combining with (9),
-$|f_{A}^{\prime}(t)|\leq\frac{C_{\mathrm{BCV}}\,\Delta}{\sqrt{(1-t)\sigma_{\textup{{\bf q}}}^{2}+t\sigma_{\textup{{\bf p}}}^{2}+1}}.$
+$|f_{A}^{\prime}(t)|\leq\frac{C_{0}\,\Delta}{\sqrt{(1-t)\sigma_{\textup{{\bf q}}}^{2}+t\sigma_{\textup{{\bf p}}}^{2}+1}}.$
 Integrating over $t$ gives
-$\bigl|f_{A}(1)-f_{A}(0)\bigr|\leq C_{\mathrm{BCV}}\,\Delta\int_{0}^{1}\frac{\mathrm{d}t}{\sqrt{(1-t)(\sigma_{\textup{{\bf q}}}^{2}+1)+t(\sigma_{\textup{{\bf p}}}^{2}+1)}}.$
+$\bigl|f_{A}(1)-f_{A}(0)\bigr|\leq C_{0}\,\Delta\int_{0}^{1}\frac{\mathrm{d}t}{\sqrt{(1-t)(\sigma_{\textup{{\bf q}}}^{2}+1)+t(\sigma_{\textup{{\bf p}}}^{2}+1)}}.$
 Write $(1-t)(\sigma_{\textup{{\bf q}}}^{2}+1)+t(\sigma_{\textup{{\bf p}}}^{2}+1)=b+at$ with $b=\sigma_{\textup{{\bf q}}}^{2}+1$ and $a=\sigma_{\textup{{\bf p}}}^{2}-\sigma_{\textup{{\bf q}}}^{2}$.
@@
 Thus, for every event $A$,
-$\bigl|\mathbb{P}(S_{\textup{{\bf p}}}\in A)-\mathbb{P}(S_{\textup{{\bf q}}}\in A)\bigr|\leq\frac{2C_{\mathrm{BCV}}\,\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+\sqrt{\sigma_{\textup{{\bf q}}}^{2}+1}}.$
-Taking $\sup_{A}$ proves (2.1). ∎
-2.2 Proof of Theorem 1.2
-We make use of a concentration-variance inequality [13, Disp.(18)], which immediately implies the following “pigeonhole” lemma: [^4] We thank M.
-Lutz [12] for bringing this to our attention and improving our original lower bound by a factor of $4/3$.
+$\bigl|\mathbb{P}(S_{\textup{{\bf p}}}\in A)-\mathbb{P}(S_{\textup{{\bf q}}}\in A)\bigr|\leq\frac{2C_{0}\,\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+\sqrt{\sigma_{\textup{{\bf q}}}^{2}+1}}.$
+Taking $\sup_{A}$ proves the claim. ∎
+2.2. Proof of Theorem 1.2
+We make use of a concentration-variance inequality stated in [15, Disp.
+(18)] and proved therein, which immediately implies the following second-moment extraction lemma.
+We thank L.
+Mattner [14] for bringing this to our attention and improving our original lower bound by a factor of $4/3$.
+We also thank him for the attributional note, “This is due to Paul Lévy in a sharper version, and was in essentially the present version apparently independently discovered by [5, Corollary 2.2]”.
 Lemma 2.2.
@@
 $\sup_{k\in\mathbb{Z}}g(k)\;\geq\;\frac{G^{3/2}}{4\sqrt{J}+\sqrt{G}}.$
-Proof (sketch).
-Normalize $g$ to be a probability measure $P$ via $P(\left\{k\right\})=g(k)/G$ and let $X$ be distributed according to $P$. Using the variational characterization of variance, $Var(X)\leq J/G$.
-By Mattner–Schulz [13, Disp. (18)], for any law with variance $\sigma^{2}$ and any $h>0$,
+Proof.
+Normalize $g$ to be a probability measure $P$ via $P(\left\{k\right\})=g(k)/G$ and let $X$ be distributed according to $P$; then $Var(X)\leq J/G$.
+By Lévy’s extraction principle, [15, Disp.
+(18)], for any law with variance $\sigma^{2}$ and any $h>0$,
 $\sup_{x\in\mathbb{R}}P\bigl((x,x+h)\bigr)\;\geq\;\frac{h}{\sqrt{h^{2}+12\sigma^{2}}}.$
-Apply this with $h=1$ and $\sigma^{2}=\operatorname{Var}(X)$. Since $P$ is supported on $\mathbb{Z}$, any interval of length $1$ contains at most one atom, and hence
+Apply this with $h=1$ and $\sigma^{2}=\operatorname{Var}(X)$. Since $P$ is supported on $\mathbb{Z}$,
 $\sup_{x\in\mathbb{R}}P((x,x+1))=\max_{k\in\mathbb{Z}}P(\{k\})=\frac{1}{G}\max_{k\in\mathbb{Z}}g(k).$
 Therefore,
-$\frac{1}{G}\max_{k}g(k)\;\geq\;\frac{1}{\sqrt{1+12\operatorname{Var}(X)}}.$
-Using $\operatorname{Var}(X)\leq J/G$,
-$\max_{k}g(k)\;\geq\;\frac{G}{\sqrt{1+12(J/G)}}=\frac{G^{3/2}}{\sqrt{G+12J}}.$
-Finally, since $\sqrt{G+12J}\leq\sqrt{G}+4\sqrt{J}$, we conclude
-$\max_{k\in\mathbb{Z}}g(k)\;\geq\;\frac{G^{3/2}}{\,4\sqrt{J}+\sqrt{G}\,}$,
-which is the claimed bound. ∎
+$\max_{k}g(k)$ $\geq$ $\frac{G}{\sqrt{1+12\operatorname{Var}(X)}}\geq\frac{G}{\sqrt{1+12(J/G)}}=\frac{G^{3/2}}{\sqrt{G+12J}}.$
+Finally, since $\sqrt{G+12J}\leq\sqrt{G}+4\sqrt{J}$, we conclude $\max_{k\in\mathbb{Z}}g(k)\;\geq\;\frac{G^{3/2}}{\,4\sqrt{J}+\sqrt{G}\,}$, as claimed. ∎
 As above, $(\textup{{\bf p}},\textup{{\bf q}})$ is a dominating pair with $p_{i}\geq q_{i}$ and associated Poisson binomials $S_{\textup{{\bf p}}}$, $S_{\textup{{\bf q}}}$.
@@
 To this end, we define the $G$ and $J$ functionals — both in terms of the non-negative (by the monotone coupling) measure $g$ on $\mathbb{Z}$:
-$g(k):=\mathbb{P}(X\geq k)-\mathbb{P}(Y\geq k),\qquad k\in\mathbb{Z}.$ (10)
+(10) $g(k):=\mathbb{P}(X\geq k)-\mathbb{P}(Y\geq k),\qquad k\in\mathbb{Z}.$
 Then $G=\sum_{k\in\mathbb{Z}}g(k)$ has the more familiar form
@@
 and so
-$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\;\geq\;\sup_{k\in\mathbb{Z}}g(k).$ (11)
+(11) $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\;\geq\;\sup_{k\in\mathbb{Z}}g(k).$
 Finally, $J=\sum_{k\in\mathbb{Z}}(k-\mu_{\textup{{\bf p}}})^{2}g(k)$ is defined as in Lemma 2.2. The technical core of the argument hinges on upper bounding $J$:
@@
 where $\mathbb{E}_{t}$ denotes expectation under the product law with parameters $\{r_{j}(t)\}_{j}$. Hence
-$J=\int_{0}^{1}\sum_{i=1}^{n}\Delta_{i}\,\mathbb{E}_{t}\bigl[(T_{i}(t)+1-\mu_{\textup{{\bf p}}})^{2}\bigr]\,\mathrm{d}t.$ (12)
+(12) $J=\int_{0}^{1}\sum_{i=1}^{n}\Delta_{i}\,\mathbb{E}_{t}\bigl[(T_{i}(t)+1-\mu_{\textup{{\bf p}}})^{2}\bigr]\,\mathrm{d}t.$
 We bound $\mathbb{E}_{t}[(T_{i}(t)+1-\mu_{\textup{{\bf p}}})^{2}]$ uniformly in $t,i$ via the decomposition $\mathbb{E}[(Z-\alpha)]^{2}=\operatorname{Var}(Z)+(\mathbb{E}Z-\alpha)^{2}$, where $Z=T_{i}(t)+1$ and $\alpha=\mu_{\textup{{\bf p}}}$.
@@
 Since $T_{i}(t)=S(t)-X_{i}(t)$ with $X_{i}(t)$ Bernoulli,
-$\operatorname{Var}_{t}(T_{i}(t))\leq\operatorname{Var}_{t}(S(t))\leq\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}.$ (13)
+(13) $\operatorname{Var}_{t}(T_{i}(t))\leq\operatorname{Var}_{t}(S(t))\leq\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}.$
 To bound the bias term, we have
@@
 Combining the two cases proves the claim. ∎
-2.3 Proof of Theorem 1.3
+2.3. Proof of Theorem 1.3
 Both inequalities invoke the data processing inequality; the first in the form
@@
 Given these, the claim is an immediate consequence of Theorem 1.2.
-2.4 Proof of Lemma 1.4
+2.4. Proof of Lemma 1.4
 Proof.
@@
 Moreover, for product measures one has
-$\operatorname{TV}(P\times R,\;Q\times S)\leq\operatorname{TV}(P,Q)+\operatorname{TV}(R,S),$ (14)
+(14) $\operatorname{TV}(P\times R,\;Q\times S)\leq\operatorname{TV}(P,Q)+\operatorname{TV}(R,S),$
 whence
@@
 This completes the proof. ∎
-Acknowledgements
+Appendix A Majorization kernels for complementary Bernoulli products
+Fix $n\geq 2$ and write $\Omega:=\{0,1\}^{n}$. For a parameter vector $\alpha=(\alpha_{1},\dots,\alpha_{n})\in[0,1]^{n}$ and $s\in\{+,-\}$ define the product measures
+$\mu_{\alpha}^{\,s}:=\bigotimes_{i=1}^{n}\operatorname{Ber}\!\left(\frac{1}{2}+s\,\frac{\alpha_{i}}{2}\right).$
+Thus $\mu_{\alpha}^{\,+}$ and $\mu_{\alpha}^{\,-}$ are complementary in the sense that $\mu_{\alpha}^{\,-}$ is obtained from $\mu_{\alpha}^{\,+}$ by flipping each bit.
+We show that, in this complementary setting, not only homogenization but any majorization of the parameter vector can be realized by a Markov kernel.
+We write $\beta\prec\alpha$ to denote that $\beta$ is majorized by $\alpha$, i.e., letting $\alpha^{\downarrow}$ and $\beta^{\downarrow}$ be the vectors obtained by sorting the coordinates of $\alpha$ and $\beta$ in nonincreasing order, we have
+$\sum_{i=1}^{k}\beta_{i}^{\downarrow}\ \leq\ \sum_{i=1}^{k}\alpha_{i}^{\downarrow}\quad\text{for all }k=1,\dots,n-1,\qquad\text{and}\qquad\sum_{i=1}^{n}\beta_{i}^{\downarrow}\ =\ \sum_{i=1}^{n}\alpha_{i}^{\downarrow}.$
+(see [1, Ch.
+1]).
+We thank L.
+Mattner [14] for this suggestion; our original result had proven Lemma A.1 only for $\lambda=\frac{1}{2}$ and Theorem A.2 only for homogenization via a compactness argument.
+A two–coordinate kernel for a general TT–transform
+The basic step is a Markov kernel acting on two coordinates which implements the standard TT–transform (a convex combination of the identity and a transposition).
+Lemma A.1 (Two–coordinate TT–transform kernel).
+Fix $a,b\in[0,1]$ and $\lambda\in[0,1]$, and define
+$a^{\prime}:=\lambda a+(1-\lambda)b,\qquad b^{\prime}:=\lambda b+(1-\lambda)a.$
+Let
+$\eta:=\begin{cases}\displaystyle\frac{\lambda(1-\lambda)(a-b)^{2}}{2(1-ab)},&ab<1,\\ [1ex]0,&ab=1,\end{cases}$
+and consider the row–stochastic $4\times 4$ matrix (indexed by $00,01,10,11$)
+(15) $M_{\lambda}(a,b):=\left(\begin{array}[]{cccc}1&0&0&0\\ \eta&\lambda-\eta&(1-\lambda)-\eta&\eta\\ \eta&(1-\lambda)-\eta&\lambda-\eta&\eta\\ 0&0&0&1\end{array}\right).$
+Then all entries of $M_{\lambda}(a,b)$ are nonnegative, and for both $s\in\{+,-\}$,
+$\Bigl(\operatorname{Ber}\!\left(\tfrac{1}{2}+s\,\tfrac{a}{2}\right)\otimes\operatorname{Ber}\!\left(\tfrac{1}{2}+s\,\tfrac{b}{2}\right)\Bigr)\,M_{\lambda}(a,b)=\operatorname{Ber}\!\left(\tfrac{1}{2}+s\,\tfrac{a^{\prime}}{2}\right)\otimes\operatorname{Ber}\!\left(\tfrac{1}{2}+s\,\tfrac{b^{\prime}}{2}\right).$
+Proof.
+Row–stochasticity is immediate from (15). For nonnegativity, note that for $a,b\in[0,1]$,
+$2(1-ab)-(a-b)^{2}=2-a^{2}-b^{2}\geq 0,$
+hence $(a-b)^{2}\leq 2(1-ab)$ whenever $ab<1$. Therefore $\eta\leq\lambda(1-\lambda)\leq\min\{\lambda,1-\lambda\}$, which implies $\lambda-\eta\geq 0$ and $(1-\lambda)-\eta\geq 0$.
+To verify the mapping property, write (for $s\in\{+,-\}$) $p_{s}:=\tfrac{1}{2}+s\tfrac{a}{2}$, $q_{s}:=\tfrac{1}{2}+s\tfrac{b}{2}$ and $p_{s}^{\prime}:=\tfrac{1}{2}+s\tfrac{a^{\prime}}{2}$, $q_{s}^{\prime}:=\tfrac{1}{2}+s\tfrac{b^{\prime}}{2}$.
+Let $x_{s}$ be the stochastic row vector of the input law $\operatorname{Ber}(p_{s})\otimes\operatorname{Ber}(q_{s})$ ordered as $(00,01,10,11)$, and similarly let $y_{s}$ be that of $\operatorname{Ber}(p_{s}^{\prime})\otimes\operatorname{Ber}(q_{s}^{\prime})$.
+A direct (routine) calculation using (15) shows that $x_{s}\,M_{\lambda}(a,b)=y_{s}$ for $s=+$ and for $s=$-.
+∎
+Finite decomposition via majorization
+Lemma A.1 supplies the probabilistic implementation of a single TT–transform. The following theorem is the desired global statement.
+Theorem A.2 (Majorization kernel for complementary products).
+Fix $n\geq 2$ and $\alpha,\beta\in[0,1]^{n}$ with $\beta\prec\alpha$. Then there exists a Markov kernel $K:\Omega\to\mathcal{P}(\Omega)$ such that simultaneously, for both $s\in\{+,-\}$,
+$\mu_{\alpha}^{\,s}K=\mu_{\beta}^{\,s}.$
+Proof.
+By the Muirhead–Hardy–Littlewood–Pólya lemma (see [1, Lemma B.1]), there exist indices $(i_{t},j_{t})$ and parameters $\lambda_{t}\in[0,1]$, $t=1,\dots,m$, such that if we define $\alpha^{(0)}:=\alpha$ and for each $t\geq 1$ update only coordinates $(i_{t},j_{t})$ via
+$\bigl(\alpha^{(t)}_{i_{t}},\alpha^{(t)}_{j_{t}}\bigr)=\bigl(\lambda_{t}\alpha^{(t-1)}_{i_{t}}+(1-\lambda_{t})\alpha^{(t-1)}_{j_{t}},\ \lambda_{t}\alpha^{(t-1)}_{j_{t}}+(1-\lambda_{t})\alpha^{(t-1)}_{i_{t}}\bigr)$,
+with all other coordinates unchanged, then $\alpha^{(m)}=\beta$.
+For each step $t$, let $K_{t}$ be the Markov kernel on $\Omega$ that acts only on coordinates $(i_{t},j_{t})$ as follows: given $x\in\Omega$, sample
+$(x^{\prime}_{i_{t}},x^{\prime}_{j_{t}})\sim M_{\lambda_{t}}\!\left(\alpha^{(t-1)}_{i_{t}},\alpha^{(t-1)}_{j_{t}}\right)(x_{i_{t}},x_{j_{t}})$
+(using Lemma A.1), and set $x^{\prime}_{k}=x_{k}$ for $k\neq i_{t},j_{t}$. By independence of coordinates and Lemma A.1, for both $s\in\{+,-\}$ we have
+$\mu_{\alpha^{(t-1)}}^{\,s}K_{t}=\mu_{\alpha^{(t)}}^{\,s}.$
+Finally, define the composed kernel $K:=K_{1}K_{2}\cdots K_{m}$. Iterating the last display yields $\mu_{\alpha}^{\,s}K=\mu_{\beta}^{\,s}$ for both $s\in\{+,-\}$. ∎
+Corollary A.3 (Homogenization as a special case).
+Let $\alpha\in[0,1]^{n}$ and put $\bar{\alpha}:=\frac{1}{n}\sum_{i=1}^{n}\alpha_{i}$. Then there exists a Markov kernel $K$ on $\Omega$ such that for both $s\in\{+,-\}$,
+$\mu_{\alpha}^{\,s}K=\mu_{\bar{\alpha}\mathbf{1}}^{\,s}.$
+Proof.
+It is immediate that $\bar{\alpha}\mathbf{1}\prec\alpha$. Apply Theorem A.2 with $\beta=\bar{\alpha}\mathbf{1}$. ∎
+Acknowledgments
 Some of the proofs were assisted by ChatGPT Pro.
@@
 We are grateful to Lutz Mattner and Bero Roos for detailed and insightful comments on the manuscript.
-Appendix A Homogenization is a Markov Kernel in the symmetric case
-We show that for symmetric $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ — i.e., those satisfying $\textup{{\bf p}}=\textup{{\bf 1}}-\textup{{\bf q}}$, homogenization can be realized by a Markov kernel.
-We first consider the case $n=2$:
-Lemma A.1 (averaging centered Bernoullis).
-Let $\alpha,\beta\in(0,1)$ and put $\gamma=(\alpha+\beta)/2$. Then there exists a randomized function $F_{\alpha,\beta}:\left\{0,1\right\}^{2}\to\left\{0,1\right\}^{2}$ with the following property:
-$X\sim\operatorname{Ber}\left(\frac{1}{2}+\frac{\alpha}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}+\frac{\beta}{2}\right)$ $\implies$ $F_{\alpha,\beta}(X)\sim\operatorname{Ber}\left(\frac{1}{2}+\frac{\gamma}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}+\frac{\gamma}{2}\right),$
-$Y\sim\operatorname{Ber}\left(\frac{1}{2}-\frac{\alpha}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}-\frac{\beta}{2}\right)$ $\implies$ $F_{\alpha,\beta}(Y)\sim\operatorname{Ber}\left(\frac{1}{2}-\frac{\gamma}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}-\frac{\gamma}{2}\right).$
-Proof.
-As in the proof of [9, Lemma 2.3], it suffices to write down the stochastic vectors representing $X,Y$ and $\hat{X}\sim\operatorname{Ber}\left(\frac{1}{2}+\frac{\gamma}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}+\frac{\gamma}{2}\right),\hat{Y}\sim\operatorname{Ber}\left(\frac{1}{2}-\frac{\gamma}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}-\frac{\gamma}{2}\right)$, the stochastic matrix $M$ representing $F_{\alpha,\beta}$, and to verify that $\hat{X}=XM$ and $\hat{Y}=YM$.
-For readability, we mildly abuse notation and blur the distinction between the random variable $Z\sim\operatorname{Ber}(p)\otimes\operatorname{Ber}(q)$ and its stochastic vector:
-$Z=\left(\mathbb{P}(Z=00),\mathbb{P}(Z=01),\mathbb{P}(Z=10),\mathbb{P}(Z=11)\right)$ $=$ $\left((1-p)(1-q),(1-p)q,p(1-q),pq\right).$
-With this convention, the four stochastic vectors are:
-$X$ $=$ $\left(\left(\frac{1}{2}-\frac{\alpha}{2}\right)\left(\frac{1}{2}-\frac{\beta}{2}\right),\left(\frac{1}{2}-\frac{\alpha}{2}\right)\left(\frac{1}{2}+\frac{\beta}{2}\right),\left(\frac{1}{2}+\frac{\alpha}{2}\right)\left(\frac{1}{2}-\frac{\beta}{2}\right),\left(\frac{1}{2}+\frac{\alpha}{2}\right)\left(\frac{1}{2}+\frac{\beta}{2}\right)\right),$
-$Y$ $=$ $\left(\left(\frac{1}{2}+\frac{\alpha}{2}\right)\left(\frac{1}{2}+\frac{\beta}{2}\right),\left(\frac{1}{2}+\frac{\alpha}{2}\right)\left(\frac{1}{2}-\frac{\beta}{2}\right),\left(\frac{1}{2}-\frac{\alpha}{2}\right)\left(\frac{1}{2}+\frac{\beta}{2}\right),\left(\frac{1}{2}-\frac{\alpha}{2}\right)\left(\frac{1}{2}-\frac{\beta}{2}\right)\right),$
-$\hat{X}$ $=$ $\left(\left(\frac{1}{2}-\frac{\gamma}{2}\right)\left(\frac{1}{2}-\frac{\gamma}{2}\right),\left(\frac{1}{2}-\frac{\gamma}{2}\right)\left(\frac{1}{2}+\frac{\gamma}{2}\right),\left(\frac{1}{2}+\frac{\gamma}{2}\right)\left(\frac{1}{2}-\frac{\gamma}{2}\right),\left(\frac{1}{2}+\frac{\gamma}{2}\right)\left(\frac{1}{2}+\frac{\gamma}{2}\right)\right),$
-$\hat{Y}$ $=$ $\left(\left(\frac{1}{2}+\frac{\gamma}{2}\right)\left(\frac{1}{2}+\frac{\gamma}{2}\right),\left(\frac{1}{2}+\frac{\gamma}{2}\right)\left(\frac{1}{2}-\frac{\gamma}{2}\right),\left(\frac{1}{2}-\frac{\gamma}{2}\right)\left(\frac{1}{2}+\frac{\gamma}{2}\right),\left(\frac{1}{2}-\frac{\gamma}{2}\right)\left(\frac{1}{2}-\frac{\gamma}{2}\right)\right).$
-With $\eta:=\frac{(\alpha-\beta)^{2}}{8(1-\alpha\beta)}$, we define the $4\times 4$ row-stochastic matrix $M$:
-$M=\left(\begin{array}[]{cccc}1&0&0&0\\ \eta&\frac{1}{2}-\eta&\frac{1}{2}-\eta&\eta\\ \eta&\frac{1}{2}-\eta&\frac{1}{2}-\eta&\eta\\ 0&0&0&1\end{array}\right).$ (19)
-Verifying the identities $\hat{X}=XM$ and $\hat{Y}=YM$ is tedious but quite straightforward.
-Obviously, the rows of $M$ sum to $1$.
-It only remains to show that the entries of $M$ are nonnegative, or equivalently, that $\eta=\frac{(\alpha-\beta)^{2}}{8(1-\alpha\beta)}\leq 1/2$.
-The latter is equivalent to ${\alpha^{2}-2\alpha\beta+\beta^{2}}\leq{4(1-\alpha\beta)}$.
-Expanding and collecting like terms yields the equivalent claim $4-(\alpha+\beta)^{2}\geq 0$, which obviously holds for $\alpha,\beta\in(0,1)$.
-∎
-For general $n$, the pairwise averaging kernel $M=M(\alpha_{i},\beta_{i})$ defined in (19) can be applied repeatedly to the pairs $(i,j)$.
-Although the kernel itself changes with each iteration, a standard compactness argument shows that for any symmetric pair $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ there exists a Markov kernel $M_{\textup{{\bf p}},\textup{{\bf q}}}$ simultaneously achieving $M_{\textup{{\bf p}},\textup{{\bf q}}}:\operatorname{Ber}(\textup{{\bf p}})\mapsto\operatorname{Ber}(\bar{p})^{\otimes n}$ and $M_{\textup{{\bf p}},\textup{{\bf q}}}:\operatorname{Ber}(\textup{{\bf q}})\mapsto\operatorname{Ber}(\bar{q})^{\otimes n}$.
+This research was supported in part by the Israel Science Foundation ISF grant 581/25 and the Binational Science Foundation BSF grant 2024243.